### v1

Abstract
Fast and accurate simulation of soft tissue deformation is a critical factor for surgical robotics and medical training. In this paper, we introduce a novel physics-informed neural simulator that approximates soft tissue deformations in a realistic and real-time manner. Our framework integrates Kelvinlet-based priors into neural simulators, making it the first approach to leverage Kelvinlets for residual learning and regularization in data-driven soft tissue modeling. By incorporating large-scale Finite Element Method (FEM) simulations of both linear and nonlinear soft tissue responses, our method improves neural network predictions across diverse architectures, enhancing accuracy and physical consistency while maintaining low latency for real-time performance. We demonstrate the effectiveness of our approach by performing accurate surgical maneuvers that simulate the use of standard laparoscopic tissue grasping tools with high fidelity. These results establish Kelvinlet-augmented learning as a powerful and efficient strategy for real-time, physics-aware soft tissue simulation in surgical applications.
Keywords:
Neural Simulators Soft Tissue Deformation Kelvinlet
1 Introduction
Soft tissue deformation modeling is a critical need in computational biomechanics for surgical simulation, image registration, and robotic-assisted interventions. Accurate tissue simulation supports surgical planning, image-guided predictions of soft-tissue changes, and precise haptic feedback for robotics. Traditional soft tissue deformation modeling relies on physics-based methods like Finite Element (FE) models, which offer high accuracy but are computationally expensive for real-time use [27]. Alternative models, including mass-spring systems [17], mesh-free methods [2], and particle-based methods [14], approximate tissue behavior, but require substantial computation and struggle to achieve realistic modeling.
Recent advances in deep learning have enabled efficient, data-driven approximations of soft tissue deformations. CNNs, GNNs, and Transformers [20] can learn complex deformation patterns from simulations or real-world data, enabling real-time tissue simulation. However, they face two key limitations: 1) poor generalization to unseen force distributions and boundary conditions, leading to unreliable extrapolation, and 2) the risk of producing physically implausible deformations, particularly when trained on limited datasets.
To overcome these limitations, physics-informed machine learning integrates fundamental physical laws into neural network training, offering a promising approach to improving accuracy and generalization. Physics-Informed Neural Networks (PINNs) [22] enforce differential equations via soft constraints, while hybrid methods like differentiable FEM solvers [25] and coarse-to-fine neural surrogates [12] aim to balance accuracy and efficiency. However, these approaches are computationally demanding, particularly for large soft-tissue meshes, and require careful consideration of numerical stability and differentiability.
We propose a physics-informed neural network framework that leverages Kelvinlets [6] to improve the accuracy and physical consistency of soft tissue deformation modeling. As a closed-form Green’s function solution to the Navier-Cauchy equations, Kelvinlets provide an efficient and analytically tractable prior. In this paper, we introduce two hybrid approaches: 1) using Kelvinlets as a baseline and learning the residual (modeling error) via a neural network, and 2) incorporating Kelvinlets as a regularizer for a neural network predicting the full deformation field. We generated a finite element dataset of liver deformations, using a linear model without preload and a nonlinear model with gravity-induced preload. We demonstrate that our proposed Neural-Augmented Kelvinlet outperforms purely data-driven models in generalization, physical realism, and learning efficiency, enabling real-time realistic modeling of soft-tissue deformations with multiple graspers.
Figure 1: Visualization of our Neural Augmented Simulator with realistic textures.
Contributions. Our work makes the following key contributions [^1] Repository available at https://github.com/Anon92373/Neural-Kelvinlet.:
  1. 1.
    We present a dataset of 20,800 FEM solutions for data-driven soft tissue modeling, with 10,400 linear and 10,400 nonlinear simulations. It includes 5,000 single-grasper and 5,400 multi-grasper cases.
  2. 2.
    We provide a GPU-accelerated Kelvinlet PyTorch implementation that enables efficient training and real-time inference, making physics-informed deformation modeling accessible for interactive applications.
  3. 3.
    We introduce two hybrid approaches that integrate Kelvinlets into data-driven soft tissue deformation modeling and conduct extensive experiments to demonstrate their superiority over purely physics-informed and data-driven baselines across diverse neural architectures.
2 Kelvinlets for Soft Tissue Deformation
2.1 Navier-Cauchy Equations and Green’s Function Solution
The displacement field $\mathbf{u}:\mathbb{R}^{3}\to\mathbb{R}^{3}$ of an isotropic, homogeneous elastic medium under an applied body force density $\mathbf{f}:\mathbb{R}^{3}\to\mathbb{R}^{3}$ satisfies the Navier-Cauchy equations:
$\mu\nabla^{2}\mathbf{u}+\frac{\mu}{1-2\nu}\nabla(\nabla\cdot\mathbf{u})+\mathbf{f}=0,$ (1)
where $\mu>0$ is the shear modulus, and $\nu\in(0,0.5)$ is the Poisson’s ratio, ensuring a physically consistent material response. For a concentrated force $\mathbf{f}\delta(\mathbf{x}-\mathbf{x}_{s})$ applied at a source point $\mathbf{x}_{s}\in\mathbb{R}^{3}$, the Green’s function solution, known as the Kelvinlet, provides an analytical displacement field $\mathbf{u}:\mathbb{R}^{3}\setminus\{\mathbf{x}_{s}\}\to\mathbb{R}^{3}$ following:
$\mathbf{u}(\mathbf{x})=\frac{1}{4\pi\mu}\left(\frac{(1-\nu)I}{r}+\nu\frac{(\mathbf{x}-\mathbf{x}_{s})\otimes(\mathbf{x}-\mathbf{x}_{s})}{r^{3}}\right)\mathbf{f},$ (2)
where $r=\|\mathbf{x}-\mathbf{x}_{s}\|$ denotes the Euclidean distance from the source.
The classical Kelvinlet in Eq. (2) exhibits singularities at $\mathbf{x}_{s}$. To prevent numerical instabilities, we substitute $r$ with $r_{\varepsilon}=\sqrt{\|\mathbf{x}-\mathbf{x}_{s}\|^{2}+\varepsilon^{2}}$ for $\epsilon>0$, where the spatial extent of force distribution is controlled by $\varepsilon$, ensuring the solution is well-conditioned. Finally, a re-scaled solution with respect to the local displacement boundary condition $\mathbf{u}_{s}=\mathbf{u}(\mathbf{x}_{s})$ removes the dependence on both the force $\mathbf{f}$ and the modulus $\mu$, yielding the regularized Kelvinlet formulation used in this paper,
$\mathbf{u}_{\varepsilon}(\mathbf{x})=\frac{\varepsilon}{(5-6\nu)}\left(\frac{(3-4\nu)I}{r_{\varepsilon}}+\frac{(\mathbf{x}-\mathbf{x}_{s})\otimes(\mathbf{x}-\mathbf{x}_{s})}{r_{\varepsilon}^{3}}\right)\mathbf{u}_{s}.$ (3)
For multiple interaction points, the Kelvinlet model represents the total organ deformation as $\mathbf{u}(\mathbf{x})=\sum_{i}k_{i}\mathbf{u}_{\varepsilon,i}(\mathbf{x}),$ where the coefficients $k_{i}$ are optimized to satisfy displacement boundary conditions $\mathbf{u}(\mathbf{x}_{i})=\mathbf{u}_{i}$, via:
$\operatorname*{argmin}_{\mathbf{k}}(\sum_{i}\|\sum_{j}k_{j}\mathbf{u}_{\varepsilon,j}(\mathbf{x}_{i})-\mathbf{u}_{i}\|^{2})+\lambda\sum_{i}k_{i}^{2}.$ (4)
This ensures that the contributions of individual Kelvinlet functions collectively satisfy the prescribed displacements. The regularization $\lambda=0.001$ ensures a minimum energy solution. The total Kelvinlet deformation model, including the optimization, can be performed in less than 3 ms for two simultaneously engaged graspers using native GPU acceleration in pytorch on a mesh with $\sim 10\text{k}$ nodes.
2.2 Method 1: Kelvinlet-Based Residual Learning
[…]
$\mathcal{L}_{\text{residual}}(\boldsymbol{\theta})=\mathbb{E}_{(\mathbf{x}_{s},\mathbf{u}_{s})\sim p}\left[\|\mathbf{u}_{\text{true}}(\cdot;\mathbf{x}_{s},\mathbf{u}_{s})-\tilde{\mathbf{u}}(\cdot;\mathbf{x}_{s},\mathbf{u}_{s},\boldsymbol{\theta})\|^{2}_{L_{2}(\Omega)}\right],$ (5)
where $p$ represents the distribution of all possible pointwise interactions with the tissue, and $\Omega$ denotes the tissue domain. By focusing on learning only the correction beyond the Kelvinlet solution, this approach enhances efficiency and interpretability while leveraging a known analytical model as a strong prior.
2.3 Method 2: Kelvinlets as Regularization
Alternatively, a network can be trained to directly regress $\mathbf{u}_{\text{true}}$, producing an estimate $\hat{\mathbf{u}}$. To enforce physically consistent predicted deformations by the network, we leverage Kelvinlets as a regularization prior and write:
$\mathcal{L}(\boldsymbol{\theta})=\mathbb{E}_{(\mathbf{x}_{s},\mathbf{u}_{s})\sim p}\left[\left\|\mathbf{u}_{\text{true}}-\hat{\mathbf{u}}\right\|^{2}_{L_{2}(\Omega)}\right]+\lambda_{reg}\mathbb{E}_{(\mathbf{x}^{\prime}_{s},\mathbf{u}^{\prime}_{s})\sim q}\left[\left\|\mathbf{u}_{\epsilon}-\hat{\mathbf{u}}\right\|^{2}_{L_{2}(\Omega)}\right],$ (6)
where $\lambda_{\text{reg}}>0$ is the regularization coefficient. For clarity and brevity, we omit the dependencies $(\cdot;\mathbf{x}_{s},\mathbf{u}_{s})$ and $(\cdot;\mathbf{x}_{s},\mathbf{u}_{s},\boldsymbol{\theta})$ in the notation. Note that we explicitly separate $(\mathbf{x}_{s},\mathbf{u}_{s})$ from $(\mathbf{x}_{s}^{\prime},\mathbf{u}^{\prime}_{s})$ because, in practice, the two distributions $p$ and $q$ could differ, and also while Monte Carlo integration is used for both terms, the first term requires FEM simulations to obtain $\mathbf{u}_{\text{true}}$ (which is computationally expensive), whereas the second term only involves Kelvinlet evaluations (which are much faster). We will describe the distributions $p$ and $q$ used in our experiments in the following section.
For both proposed methods, when multiple graspers interact with the tissue, we employ the multi-point interaction solution of Kelvinlet and provide all interaction points along with their corresponding displacements to the network.
3 Large-Scale FEM Data
Our neural network framework trains on a large FE simulation dataset, focused on lifting liver lobes to expose the gallbladder during laparoscopic cholecystectomy. We construct an FE model from an abdominal CT image and its segmentation masks, converting them into a 3D linear tetrahedral mesh with 10,400 nodes and 44,045 tetrahedra using ‘gmsh’ [9].
Interactions of the liver and gallbladder with surrounding organs are modeled by prescribing boundary conditions. Those boundary conditions restrict regions of the liver’s exposed surface that are considered in the far field from the interaction location, as well as the distal end of the cystic ducts, arteries, and veins. FE simulations are solved using the open source getFEM package [24].
[…]
$\boldsymbol{E}=\frac{1}{2}\left(\nabla\mathbf{u}+\nabla\mathbf{u}^{T}+\nabla\mathbf{u}^{T}\nabla\mathbf{u}\right).$ (7)
In addition, this equation relates to the Cauchy-Green deformation tensor $\boldsymbol{C}$ by: $\boldsymbol{C}=2\boldsymbol{E}-\mathbf{I}$, where $\mathbf{I}$ is the identity tensor.
For our nonlinear simulations, we use a two-parameter Mooney-Rivlin model, where the strain energy density can be written as functions of the Cauchy-Green deformation tensor invariants ($\boldsymbol{I_{i}(C)}$, with $i=1,2,3$) [10]:
$\boldsymbol{W}=C_{01}\left(\boldsymbol{I_{1}(C)}\boldsymbol{I_{3}(C)}^{\frac{-1}{3}}-3\right)+C_{10}\left(\boldsymbol{I_{1}(C)}\boldsymbol{I_{3}(C)}^{\frac{-2}{3}}-3\right),$ (8)
where $C_{10}=1.62$ kPa and $C_{01}=1.97$ kPa are constitutive constants derived from ex-vivo porcine liver experiments [3].
For our linear simulations, we simulate an elastic response in the absence of gravitational load ($\boldsymbol{g}=0$). These solutions serve as baseline data for tissue deformation under small elastic input displacements, where the term $\nabla\mathbf{u}^{T}\nabla\mathbf{u}$ in Eq. (7) is neglected. Assuming isotropic material properties, the strain energy density $\boldsymbol{W}$ in terms of shear modulus $\mu$ and Poisson’s ratio $\nu$ expressed as:
$\boldsymbol{W}=\left(\frac{\mu\nu}{1-2\nu}\right)\mathrm{Tr}(\boldsymbol{E})^{2}+\mu\mathrm{Tr}(\boldsymbol{E}^{2}).$ (9)
In these linear simulations, we assign a Poisson’s ratio of 0.45 for all organs [4]. The shear modulus values are also set as follows: 0.03 kPa for the fascia [1], 0.34 kPa for the gallbladder [18], 0.69 kPa for the gallbladder wall and liver [11], and 1.07 kPa for the remaining organs [8].
3.0.1 Training Distribution, $p$.
Given a mesh partitioned into regions $\Omega_{i}$ informed by cholecystectomy surgery, nodes $\mathbf{x}_{s}$ are sampled from a mixture of Generalized von Mises-Fisher (vMF) distributions (i.e., radial densities on the surface mesh) centered in these regions:
$p(\mathbf{x}_{s})=\sum_{i}g(\Omega_{i})f_{\text{vMF}}(\mathbf{x}_{s}\mid\boldsymbol{\mu}_{i},\kappa_{i}),\vspace{-.1in}$ (10)
where $g(\Omega_{i})$ is the prior-informed probability of selecting $\Omega_{i}$, $\boldsymbol{\mu}_{i}$ indicates the center of $\Omega_{i}$, and $\kappa_{i}$ is the spread of the vMF, which is a function of the size of $\Omega_{i}$. The signals $u_{s}$ are independently sampled from a uniform distribution in a cuboid $\mathcal{C}$ defined by: $x\in(x_{1},x_{2}),y\in(y_{1},y_{2}),z\in(z_{1},z_{2})$, with volume: $V=(x_{2}-x_{1})(y_{2}-y_{1})(z_{2}-z_{1}).$ This approach ensures that node sampling is spatially structured according to surgical priors while allowing displacements to be isotropic in a bounded range.
For single grasping, ten regions $\Omega_{i}$ around the liver edge were selected, each with 250 samples. For multiple graspers, six pairwise combinations of five regions $\Omega_{i}$ were considered, with 30 samples per region. The displacement ranges are $x\in(x_{1},x_{2})=(-90,50)$ mm, $y\in(y_{1},y_{2})=(-90,50)$ mm, and $z\in(z_{1},z_{2})=(-60,90)$ mm, corresponding to the cube domain.
3.0.2 Distribution for Multiple Graspers.
For multiple interactions (e.g., two graspers), we define the joint distribution as $p(\mathbf{x}_{s}^{1},\mathbf{x}_{s}^{2})=p(\mathbf{x}_{s}^{1})p(\mathbf{x}_{s}^{2}\mid\mathbf{x}_{s}^{1})$, where $p(\mathbf{x}_{s}^{1})$ follows Eq. (10). To ensure that $\mathbf{x}_{s}^{2}$ is not sampled from the same region $\Omega_{k}$ as $\mathbf{x}_{s}^{1}$, we modify the region probabilities $g(\Omega_{i})$ by setting $g(\Omega_{k})=0$ and adjusting $g(\Omega_{i\neq k})$ based on surgical priors, and then sampling $\mathbf{x}_{s}^{2}$ from $p(\mathbf{x}_{s}^{2}\mid\mathbf{x}_{s}^{1})$.
3.0.3 Regularization Distribution, $q$.
For Kelvinlet regularization, interaction points are sampled uniformly from the surface nodes. Initial displacements are then drawn uniformly from a cone centered around the surface normal at $\mathbf{x}_{s}$, with a maximum opening angle of $\alpha_{\max}$.
4 Texture Rendering for Realistic Simulations
Our visualization pipeline combines real-time tissue deformation and high-fidelity rendering to create a realistic simulation (see Figure 1 and the supplementary videos). In a preprocessing stage, surface meshes are extracted from the segmented high-resolution volumetric mesh, assigned texture coordinates, and enriched with procedurally generated textures using Worley noise [26]. Interpolation weights are computed to map low-resolution deformations to the high-resolution mesh [16], optimized using a Bounding Volume Hierarchy (BVH) [5], and UV coordinates are transferred to the volumetric mesh. At runtime, the system dynamically updates surface meshes to accommodate added or removed tetrahedral elements, interpolates high-resolution vertex positions based on low-resolution deformations, smooths exposed internal structures, and recalculates surface normals. Rendering employs pre-integrated diffuse shading [19], triplanar texture projection for inner tissue surfaces [7], and post-processing effects within Unity to achieve realistic tissue appearance. The pipeline achieves near-real-time performance while preserving visual fidelity during dynamic deformations.
5 Experiments
5.0.1 Data.
Our dataset consists of large-scale FEM simulations of soft tissue deformation. Each node in the mesh has a feature vector: $\mathbf{f}=\left[\mathbf{x};\mathbf{u}_{s};a\right]\in\mathbb{R}^{7}$. Where $\mathbf{x}\in\mathbb{R}^{3}$ is the undeformed position, $\mathbf{u}_{s}\in\mathbb{R}^{3}$ is the initial displacement, and $a\in\{0,1\}$ indicates whether the node is actively influenced by external forces. This distinction helps differentiate constrained nodes from passively deforming ones. We create a dataset of 20,800 FEM solutions with 10,400 linear elastic simulations and 10,400 nonlinear material responses. Of these, 5,000 correspond to single manipulations and 5,400 to multiple grasper manipulations.
5.0.2 Neural Architectures.
We evaluate Kelvinlet-based priors using four neural architectures: U-Net [15], SAB [13] (set attention), GraphGPS [23] (a Graph Neural Network and Transformer hybrid), and PointNet [21] (a universal function approximator for sets). We modify PointNet and GraphGPS by reducing layers, U-Net by replacing convolutions with shared MLPs to form a DeepSet, and SAB by using two layers with an MLP regressor head.
Each model is trained separately on two datasets: one with individual graspers and another with multiple graspers interacting. For each dataset, models are trained under three settings: (1) a baseline without physics priors (i.e., evaluating solely the network), (2) residual learning predicting deviations from Kelvinlets, and (3) physics-informed regularization enforcing Kelvinlet constraints.
5.0.3 Hyperparameters.
We set shear modulus in Kelvinlet to $\mu=0.72$ kPa, Poisson’s ratio $\nu=0.45$, and regularization parameter $\epsilon=0.05$, to reflect soft tissue properties and ensure numerical stability. For linear and nonlinear simulations we use $\lambda_{\text{reg}}=1$ and $\lambda_{\text{reg}}=0.1$, respectively, to emphasize linearity of Kelvinlet.
5.1 Results and Analysis
Figure 2: Qualitative comparison of deformation predictions for the best (PointNet) and worst (GraphGPS) models in linear (top) and nonlinear (bottom) regimes.
To quantitatively assess model performance, we introduce the Deformation Capture Mean (DCM) metric, which measures how well a predicted deformation field approximates the ground-truth displacement field. The DCM is computed as:
$\text{DCM}=100\times\left(1-\mathbb{E}_{(\mathbf{x}_{s},\mathbf{u}_{s})\sim p}\left[\frac{\left\|\mathbf{u}_{\text{true}}-\hat{\mathbf{u}}\right\|_{L_{2}(\Omega)}}{\|\mathbf{u}_{\text{true}}\|_{L_{2}(\Omega)}}\right]\right).$ (11)
Tables 1 and 2 report DCMs across models and settings. Each experiment runs with three random seeds, reporting mean and standard deviation. Models perform better on linear cases, as nonlinear deformations introduce gravity effects, non-proportional stress-strain relationships, and greater sensitivity to perturbations. Notably, Kelvinlet priors consistently improve DCMs, despite being derived for linear elasticity, showing their effectiveness even in nonlinear regimes.
Table 1: DCM of models with/without Kelvinlet priors in linear simulations.
Model Indiv. Graspers Comb. Graspers
Base +Kelv. Res. +Kelv. Reg. Base +Kelv. Res. +Kelv. Reg.
U-Net[15] 92.11 (0.12) 93.56 (0.10) 94.04 (0.11) 88.23 (0.08) 89.67 (0.09) 90.01 (0.08)
GraphGPS[23] 88.60 (0.21) 89.45 (0.15) 89.61 (0.23) 87.33 (0.06) 87.92 (0.02) 88.01 (0.04)
PointNet[21] 94.23 (0.11) 94.66 (0.09) 96.54 (0.09) 89.79 (0.07) 90.11 (0.07) 91.22 (0.07)
SAB[13] 93.42 (0.11) 94.60 (0.04) 96.03 (0.04) 88.05 (0.09) 89.91 (011) 91.02 (0.04)
Kelvinlet 65.24 (0) 64.33 (0)
Table 2: DCM of models with/without Kelvinlet priors in nonlinear simulations.
Model Indiv. Graspers Comb. Graspers
Base +Kelv. Res. +Kelv. Reg. Base +Kelv. Res. +Kelv. Reg.
U-Net[15] 80.56 (0.10) 84.22 (0.09) 84.88 (0.11) 76.91 (0.05) 79.83 (0.08) 82.10 (0.21)
GraphGPS[23] 77.52 (0.17) 80.63 (0.13) 82.11 (0.09) 73.72 (0.15) 77.92 (0.10) 80.66 (0.16)
PointNet[21] 81.25 (0.22) 84.65 (0.19) 85.05 (0.15) 79.56 (0.18) 82.66 (0.16) 83.33 (0.11)
SAB[13] 80.75 (0.19) 84.44 (0.07) 84.75 (0.10) 78.88 (0.11) 81.50 (0.15) 82.83 (0.16)
Kelvinlet 58.88 (0) 51.67 (0)
Figure 3 illustrates the trade-off between inference time and accuracy for different models across linear and nonlinear simulations. The results indicate that Kelvinlet regularization consistently improves accuracy compared to the base models, as evidenced by the upward shift in solid markers relative to their transparent counterparts. However, this enhancement in accuracy does not increase computational cost, maintaining inference times within a comparable range. Additionally, models trained on multi-grasper data (black-edged markers) exhibit similar trends to their single-grasper counterparts, suggesting that Kelvinlet regularization generalizes well across different grasping conditions. All computations were performed on a single NVIDIA A6000 GPU, ensuring a consistent hardware environment for runtime measurements.
Figure 3: Log-scaled inference time vs. accuracy (DCM%) for various models in linear and nonlinear simulations. Transparent markers denote base models, while solid markers indicate Kelvinlet-regularized versions. Black-edged markers represent multi-grasper cases, while edge-free markers denote single-grasper cases. FEM and Kelvinlet baselines are included for reference.
6 Conclusion
We introduce Neural-Augmented Kelvinlets, a physics-informed approach for real-time soft tissue deformation modeling that enhances neural network predictions in accuracy, physical consistency, and generalization while maintaining real-time performance. We utilize Kelvinlets in two settings: as a residual learning framework to refine predictions beyond an analytical baseline and as a regularization term to enforce physically plausible deformations. Large-scale FEM simulations provide a benchmark for evaluating data-driven deformation models, demonstrating that Kelvinlet integration significantly improves performance, particularly in complex multi-grasper scenarios. Our GPU-accelerated PyTorch implementation enables real-time inference, making it well-suited for surgical robotics and medical training, while a high-fidelity rendering pipeline ensures realistic soft tissue simulations in real time.
Acknowledgment
This work was partially supported by the NSF CAREER Award #2339898 and the Wellcome Leap ‘Surgery: Assess/Validate/Expand (SAVE)’ program.

### v2

@@
-
+Neural-Augmented Kelvinlet for Real-Time Soft Tissue Deformation Modeling
 Abstract
-Fast and accurate simulation of soft tissue deformation is a critical factor for surgical robotics and medical training.
-In this paper, we introduce a novel physics-informed neural simulator that approximates soft tissue deformations in a realistic and real-time manner.
-Our framework integrates Kelvinlet-based priors into neural simulators, making it the first approach to leverage Kelvinlets for residual learning and regularization in data-driven soft tissue modeling.
-By incorporating large-scale Finite Element Method (FEM) simulations of both linear and nonlinear soft tissue responses, our method improves neural network predictions across diverse architectures, enhancing accuracy and physical consistency while maintaining low latency for real-time performance.
-We demonstrate the effectiveness of our approach by performing accurate surgical maneuvers that simulate the use of standard laparoscopic tissue grasping tools with high fidelity.
-These results establish Kelvinlet-augmented learning as a powerful and efficient strategy for real-time, physics-aware soft tissue simulation in surgical applications.
-Keywords:
-Neural Simulators Soft Tissue Deformation Kelvinlet
-1 Introduction
-Soft tissue deformation modeling is a critical need in computational biomechanics for surgical simulation, image registration, and robotic-assisted interventions.
-Accurate tissue simulation supports surgical planning, image-guided predictions of soft-tissue changes, and precise haptic feedback for robotics.
-Traditional soft tissue deformation modeling relies on physics-based methods like Finite Element (FE) models, which offer high accuracy but are computationally expensive for real-time use [27].
-Alternative models, including mass-spring systems [17], mesh-free methods [2], and particle-based methods [14], approximate tissue behavior, but require substantial computation and struggle to achieve realistic modeling.
-Recent advances in deep learning have enabled efficient, data-driven approximations of soft tissue deformations.
-CNNs, GNNs, and Transformers [20] can learn complex deformation patterns from simulations or real-world data, enabling real-time tissue simulation.
-However, they face two key limitations: 1) poor generalization to unseen force distributions and boundary conditions, leading to unreliable extrapolation, and 2) the risk of producing physically implausible deformations, particularly when trained on limited datasets.
-To overcome these limitations, physics-informed machine learning integrates fundamental physical laws into neural network training, offering a promising approach to improving accuracy and generalization.
-Physics-Informed Neural Networks (PINNs) [22] enforce differential equations via soft constraints, while hybrid methods like differentiable FEM solvers [25] and coarse-to-fine neural surrogates [12] aim to balance accuracy and efficiency.
-However, these approaches are computationally demanding, particularly for large soft-tissue meshes, and require careful consideration of numerical stability and differentiability.
-We propose a physics-informed neural network framework that leverages Kelvinlets [6] to improve the accuracy and physical consistency of soft tissue deformation modeling.
-As a closed-form Green’s function solution to the Navier-Cauchy equations, Kelvinlets provide an efficient and analytically tractable prior.
-In this paper, we introduce two hybrid approaches: 1) using Kelvinlets as a baseline and learning the residual (modeling error) via a neural network, and 2) incorporating Kelvinlets as a regularizer for a neural network predicting the full deformation field.
-We generated a finite element dataset of liver deformations, using a linear model without preload and a nonlinear model with gravity-induced preload.
-We demonstrate that our proposed Neural-Augmented Kelvinlet outperforms purely data-driven models in generalization, physical realism, and learning efficiency, enabling real-time realistic modeling of soft-tissue deformations with multiple graspers.
-Figure 1: Visualization of our Neural Augmented Simulator with realistic textures.
-Contributions. Our work makes the following key contributions [^1] Repository available at https://github.com/Anon92373/Neural-Kelvinlet.:
+Accurate and efficient modeling of soft-tissue interactions is fundamental for advancing surgical simulation, surgical robotics, and model-based surgical automation.
+To achieve real-time latency, classical Finite Element Method (FEM) solvers are often replaced with neural approximations; however, naively training such models in a fully data-driven manner without incorporating physical priors frequently leads to poor generalization and physically implausible predictions.
+We present a novel physics-informed neural simulation framework that enables real-time prediction of soft-tissue deformations under complex single- and multi-grasper interactions.
+Our approach integrates Kelvinlet-based analytical priors with large-scale FEM data, capturing both linear and nonlinear tissue responses.
+This hybrid design improves predictive accuracy and physical plausibility across diverse neural architectures while maintaining the low-latency performance required for interactive applications.
+We validate our method on challenging surgical manipulation tasks involving standard laparoscopic grasping tools, demonstrating substantial improvements in deformation fidelity and temporal stability over existing baselines.
+These results establish Kelvinlet-augmented learning as a principled and computationally efficient paradigm for real-time, physics-aware soft-tissue simulation in surgical AI.
+Our code and data is available at: https://github.com/Anon92373/Neural-Kelvinlet.
+Soft tissue deformation modeling is essential in computational biomechanics for surgical simulation, image registration, and robotic-assisted interventions.
+Accurate tissue simulation enables surgical planning, image-guided predictions of soft-tissue changes, and precise haptic feedback for robotics.
+Traditional approaches rely on physics-based methods grounded in continuum mechanics, such as the Finite Element Method (FEM), Boundary Element Method (BEM), and Tensor Mass Model (TMM).
+While highly accurate and physically realistic, these methods are often computationally prohibitive for real-time applications [2, 34].
+To reduce computational cost, approximate methods such as mass-spring systems [23], mesh-free methods [3], and particle-based methods [21] are frequently used.
+However, these approximations often sacrifice realism, creating a significant simulation-to-reality gap that limits their practicality in high-fidelity applications.
+Recent advances in deep learning and neural operators have enabled fully data-driven methods that approximate soft tissue deformations with a single forward pass of a learned model.
+Convolutional neural networks (CNNs), graph neural networks (GNNs), and Transformers [26] have demonstrated the ability to capture complex deformation patterns from simulated or real-world data, enabling real-time tissue simulation.
+Despite these advances, fully data-driven approaches face two fundamental limitations: (a) limited generalization to unseen force distributions and boundary conditions, which results in unreliable extrapolation, and (b) a tendency to produce physically inconsistent deformations, particularly when trained on datasets of limited size or diversity.
+These shortcomings highlight the need for hybrid approaches that combine the expressiveness of neural models with the robustness of physics-based priors.
+To address these challenges, considerable efforts have focused on integrating physics priors into data-driven machine learning models [17].
+Broadly, these methods make learning algorithms physics-informed by introducing: (a) observational biases, such as data augmentations that embed physical constraints; (b) inductive biases in the model architecture, for example, using group-invariant or equivariant neural networks [5]; and (c) learning biases through the design of loss functions that explicitly enforce physical laws.
+Despite this progress, achieving the necessary low latency for real-time simulation while maintaining a high-performing physics-informed model with appropriately chosen biases remains a challenging and largely open problem.
+In this paper, we model soft tissue using a volumetric mesh and address the problem of predicting the displacement of the entire mesh given the observed displacements at a limited set of interaction points on its surface.
+This is analogous to determining how the entire soft tissue will deform when graspers engage multiple surface locations and are moved to specified target positions.
+While several prior studies have explored similar setup and demonstrated that neural networks can approximate the solutions produced by high-fidelity physics-based solvers such as FEM [27], these approaches often lack strong physical priors and can suffer from reduced generalization and physical inconsistency.
+To overcome these limitations, we propose a physics-informed neural framework that integrates analytical elastic priors to enhance both accuracy and physical plausibility.
+We introduce a physics-informed neural framework that leverages Kelvinlets [8], i.e., the closed-form Green’s function solutions to the Navier–Cauchy equations, as an analytically tractable and computationally efficient prior.
+Kelvinlets provide a fast, low-order approximation of elastic deformation that captures the fundamental physics of linear elasticity.
+We build on this prior through two complementary hybrid strategies: (a) residual learning: using Kelvinlets to generate a baseline displacement field and training a neural network to model the residual error; and (b) physics regularization: incorporating Kelvinlets as a soft constraint within the learning objective of a neural network that predicts the full deformation field.
+By explicitly coupling the expressiveness of neural networks with the efficiency and physical grounding of Kelvinlets, our approach achieves improved accuracy, generalization, and physical plausibility for soft tissue deformation modeling.
+To establish a unified testbed for evaluating our approach alongside prior methods, we generated a finite element dataset of liver deformations using both a linear model without preload and a nonlinear model with gravity-induced preload.
+We show that our proposed Neural-Augmented Kelvinlet framework surpasses purely data-driven models in generalization, physical realism, and learning efficiency, ultimately enabling real-time, high-fidelity modeling of soft tissue deformations involving multiple graspers.
+Contributions. Our key contributions are:
   1. 1.
@@
     We introduce two hybrid approaches that integrate Kelvinlets into data-driven soft tissue deformation modeling and conduct extensive experiments to demonstrate their superiority over purely physics-informed and data-driven baselines across diverse neural architectures.
-2 Kelvinlets for Soft Tissue Deformation
-2.1 Navier-Cauchy Equations and Green’s Function Solution
-The displacement field $\mathbf{u}:\mathbb{R}^{3}\to\mathbb{R}^{3}$ of an isotropic, homogeneous elastic medium under an applied body force density $\mathbf{f}:\mathbb{R}^{3}\to\mathbb{R}^{3}$ satisfies the Navier-Cauchy equations:
-$\mu\nabla^{2}\mathbf{u}+\frac{\mu}{1-2\nu}\nabla(\nabla\cdot\mathbf{u})+\mathbf{f}=0,$ (1)
-where $\mu>0$ is the shear modulus, and $\nu\in(0,0.5)$ is the Poisson’s ratio, ensuring a physically consistent material response.
-For a concentrated force $\mathbf{f}\delta(\mathbf{x}-\mathbf{x}_{s})$ applied at a source point $\mathbf{x}_{s}\in\mathbb{R}^{3}$, the Green’s function solution, known as the Kelvinlet, provides an analytical displacement field $\mathbf{u}:\mathbb{R}^{3}\setminus\{\mathbf{x}_{s}\}\to\mathbb{R}^{3}$ following:
-$\mathbf{u}(\mathbf{x})=\frac{1}{4\pi\mu}\left(\frac{(1-\nu)I}{r}+\nu\frac{(\mathbf{x}-\mathbf{x}_{s})\otimes(\mathbf{x}-\mathbf{x}_{s})}{r^{3}}\right)\mathbf{f},$ (2)
-where $r=\|\mathbf{x}-\mathbf{x}_{s}\|$ denotes the Euclidean distance from the source.
-The classical Kelvinlet in Eq.
-(2) exhibits singularities at $\mathbf{x}_{s}$.
-To prevent numerical instabilities, we substitute $r$ with $r_{\varepsilon}=\sqrt{\|\mathbf{x}-\mathbf{x}_{s}\|^{2}+\varepsilon^{2}}$ for $\epsilon>0$, where the spatial extent of force distribution is controlled by $\varepsilon$, ensuring the solution is well-conditioned.
-Finally, a re-scaled solution with respect to the local displacement boundary condition $\mathbf{u}_{s}=\mathbf{u}(\mathbf{x}_{s})$ removes the dependence on both the force $\mathbf{f}$ and the modulus $\mu$, yielding the regularized Kelvinlet formulation used in this paper,
-$\mathbf{u}_{\varepsilon}(\mathbf{x})=\frac{\varepsilon}{(5-6\nu)}\left(\frac{(3-4\nu)I}{r_{\varepsilon}}+\frac{(\mathbf{x}-\mathbf{x}_{s})\otimes(\mathbf{x}-\mathbf{x}_{s})}{r_{\varepsilon}^{3}}\right)\mathbf{u}_{s}.$ (3)
-For multiple interaction points, the Kelvinlet model represents the total organ deformation as $\mathbf{u}(\mathbf{x})=\sum_{i}k_{i}\mathbf{u}_{\varepsilon,i}(\mathbf{x}),$ where the coefficients $k_{i}$ are optimized to satisfy displacement boundary conditions $\mathbf{u}(\mathbf{x}_{i})=\mathbf{u}_{i}$, via:
-$\operatorname*{argmin}_{\mathbf{k}}(\sum_{i}\|\sum_{j}k_{j}\mathbf{u}_{\varepsilon,j}(\mathbf{x}_{i})-\mathbf{u}_{i}\|^{2})+\lambda\sum_{i}k_{i}^{2}.$ (4)
-This ensures that the contributions of individual Kelvinlet functions collectively satisfy the prescribed displacements.
-The regularization $\lambda=0.001$ ensures a minimum energy solution.
-The total Kelvinlet deformation model, including the optimization, can be performed in less than 3 ms for two simultaneously engaged graspers using native GPU acceleration in pytorch on a mesh with $\sim 10\text{k}$ nodes.
-2.2 Method 1: Kelvinlet-Based Residual Learning
-[…]
-$\mathcal{L}_{\text{residual}}(\boldsymbol{\theta})=\mathbb{E}_{(\mathbf{x}_{s},\mathbf{u}_{s})\sim p}\left[\|\mathbf{u}_{\text{true}}(\cdot;\mathbf{x}_{s},\mathbf{u}_{s})-\tilde{\mathbf{u}}(\cdot;\mathbf{x}_{s},\mathbf{u}_{s},\boldsymbol{\theta})\|^{2}_{L_{2}(\Omega)}\right],$ (5)
-where $p$ represents the distribution of all possible pointwise interactions with the tissue, and $\Omega$ denotes the tissue domain.
-By focusing on learning only the correction beyond the Kelvinlet solution, this approach enhances efficiency and interpretability while leveraging a known analytical model as a strong prior.
-2.3 Method 2: Kelvinlets as Regularization
-Alternatively, a network can be trained to directly regress $\mathbf{u}_{\text{true}}$, producing an estimate $\hat{\mathbf{u}}$.
-To enforce physically consistent predicted deformations by the network, we leverage Kelvinlets as a regularization prior and write:
-$\mathcal{L}(\boldsymbol{\theta})=\mathbb{E}_{(\mathbf{x}_{s},\mathbf{u}_{s})\sim p}\left[\left\|\mathbf{u}_{\text{true}}-\hat{\mathbf{u}}\right\|^{2}_{L_{2}(\Omega)}\right]+\lambda_{reg}\mathbb{E}_{(\mathbf{x}^{\prime}_{s},\mathbf{u}^{\prime}_{s})\sim q}\left[\left\|\mathbf{u}_{\epsilon}-\hat{\mathbf{u}}\right\|^{2}_{L_{2}(\Omega)}\right],$ (6)
-where $\lambda_{\text{reg}}>0$ is the regularization coefficient.
-For clarity and brevity, we omit the dependencies $(\cdot;\mathbf{x}_{s},\mathbf{u}_{s})$ and $(\cdot;\mathbf{x}_{s},\mathbf{u}_{s},\boldsymbol{\theta})$ in the notation.
-Note that we explicitly separate $(\mathbf{x}_{s},\mathbf{u}_{s})$ from $(\mathbf{x}_{s}^{\prime},\mathbf{u}^{\prime}_{s})$ because, in practice, the two distributions $p$ and $q$ could differ, and also while Monte Carlo integration is used for both terms, the first term requires FEM simulations to obtain $\mathbf{u}_{\text{true}}$ (which is computationally expensive), whereas the second term only involves Kelvinlet evaluations (which are much faster).
-We will describe the distributions $p$ and $q$ used in our experiments in the following section.
-For both proposed methods, when multiple graspers interact with the tissue, we employ the multi-point interaction solution of Kelvinlet and provide all interaction points along with their corresponding displacements to the network.
+1 Related Work
+Classical approaches based on the Finite Element Method (FEM) provide high-fidelity simulations of soft tissue deformation but are often computationally prohibitive for real-time surgical applications.
+To address this limitation, model reduction techniques such as Proper Orthogonal Decomposition (POD) have been proposed to accelerate FEM without sacrificing physical realism.
+For example, Lau et al.
+[20] employ a reduced-order linear FEM with modal warping to achieve sub-15 ms inference for liver deformation, enabling intraoperative use.
+Complementarily, Hu and Desai [14] characterize nonlinear liver material behavior through indentation experiments and validate Local Effective Moduli (LEM) estimates using hyperelastic FEM models, laying foundational work for realistic soft-tissue mechanics under large strains.
+More recently, physics-informed neural networks (PINNs) have emerged as a data-efficient alternative that integrates governing equations such as elasticity or dynamics directly into the learning objective.
+Hu et al.
+[15] introduce a PINN that encodes static linear elasticity to predict 3D stress and strain fields of liver tissue under surgical tool interaction, achieving millisecond-scale inference while retaining physical plausibility.
+Extending this paradigm, Nguyen-Le et al.
+[24] propose a PINODE framework that augments neural dynamics with mass-spring priors, enabling extrapolation from sparse motion data.
+These methods improve interpretability and generalization but often require complex training and remain constrained by solver latency or limited scalability.
+In contrast, purely data-driven surrogate models have been proposed to directly regress deformation fields from control variables.
+Ke et al.
+[6] train a feedforward network to predict liver surface displacements conditioned on retractor pose, achieving sub-millisecond inference.
+While fast and simple, such models lack physical constraints, which can reduce reliability and robustness in unseen conditions.
+On the learning-based front, mesh-based models such as PhysGNN leverage FEM supervision to train graph neural networks for real-time deformation prediction in neurosurgery [33].
+PEGNN [32] extends this with contact-aware hybrid graph reasoning, enabling accurate predictions across surgical and manipulation scenarios.
+Meanwhile, SeeSaw [9] employs stereo video and self-supervised learning to estimate sparse 3D deformations using a GNN encoder.
+Zhao et al.
+[12] introduce a diffeomorphic mapping framework that enforces cycle consistency, capturing invertible tissue motion.
+These approaches show strong empirical performance but may suffer from limited interpretability or reliance on dense supervision.
+Neural architectures such as U-Net [22], Set Attention Block (SAB) [19], and GraphGPS [29] offer modular and scalable backbones for modeling deformation fields across grid-based, set-based, and graph-structured data.
+While effective in capturing spatial patterns and long-range dependencies, these models are generally agnostic to the underlying physics of soft tissue behavior and typically lack inductive biases grounded in biomechanics, such as force equilibrium, boundary conditions, or constitutive material priors.
+Our work builds upon these foundations by integrating physically motivated priors, specifically Kelvinlet-based elastic solutions, into neural architectures through both residual learning and regularization.
+By embedding closed-form Green’s function responses from linear elasticity into the training process, our method encourages physically consistent deformation predictions even in data-sparse regimes.
+This hybrid approach improves generalization, improves physical plausibility, and remains compatible with a range of backbone architectures, while preserving the low latency performance required for real-time surgical simulation and interactive applications.
 2 Large-Scale FEM Data
-Our neural network framework trains on a large FE simulation dataset, focused on lifting liver lobes to expose the gallbladder during laparoscopic cholecystectomy.
-We construct an FE model from an abdominal CT image and its segmentation masks, converting them into a 3D linear tetrahedral mesh with 10,400 nodes and 44,045 tetrahedra using ‘gmsh’ [9].
-Interactions of the liver and gallbladder with surrounding organs are modeled by prescribing boundary conditions.
-Those boundary conditions restrict regions of the liver’s exposed surface that are considered in the far field from the interaction location, as well as the distal end of the cystic ducts, arteries, and veins.
-FE simulations are solved using the open source getFEM package [24].
-[…]
+Figure 1: Overview of the large-scale FEM data generation pipeline.
+A patient’s abdominal CT scan is first anatomically segmented (left), followed by the generation of a 3D linear tetrahedral mesh (middle).
+Diverse interactions are then applied to the mesh, and FEM simulations are performed to produce training data for ML models (right).
+The right panel shows the norms of the input (interactions) and the resulting output displacements (predicted global effects).
+Neural network–based solutions require training on large-scale finite element (FE) simulation datasets.
+In this work, we focus on laparoscopic cholecystectomy and construct an FE model from an abdominal CT image and its segmentation masks, converting them into a 3D linear tetrahedral mesh with 10,400 nodes and 44,045 tetrahedra using gmsh [11].
+As part of our contributions, we release this large-scale dataset; the details of its creation are provided below.
+The bare area of the liver, where it contacts the inferior vena cava and diaphragm, and the distal ends of the ducts and veins are fixed (non-deformable).
+Tool interaction regions are selected to simulate the initial surgical step of lifting the liver lobes to expose the gallbladder and cystic ducts.
+These regions are identified by clustering nodes along the liver ridge, based on surgical video analysis and input from our surgical collaborator.
+FE simulations are performed using the open-source getFEM package [30], applying tool-imposed displacements at those predefined clustered regions.
+This pipeline is illustrated in Figure 1.
+In our FE model, nodes with restricted motion or tool-imposed displacements are flagged.
+These boundary conditions are encoded as binary features in the neural network (1 for constrained or controlled, 0 otherwise), allowing the model to distinguish between free and constrained regions.
+Assuming a quasi-static environment, our solutions satisfy equilibrium by ensuring that the total potential energy is not only stationary but is also a minimum for permissible displacements, and is given by: $\int_{\Omega}\frac{\boldsymbol{W}(\boldsymbol{E})}{\partial\boldsymbol{\boldsymbol{E}}}:\delta\boldsymbol{E}\,dV+\int_{\Omega}\rho\boldsymbol{g}\cdot\delta\mathbf{u}\,dV=0$, where $\boldsymbol{W}(\boldsymbol{E})$ is the strain energy density, $\Omega$ is the volume of the model, $\delta\mathbf{u}$ is the variation of the displacement field, $\delta\boldsymbol{E}$ is the variation of the Green-Lagrange strain tensor, $\rho$ is the density of the material (1000 $kg/m^{3}$), and $\boldsymbol{g}$ is the acceleration due to gravity that points downward along the y-axis (9.81 $m/s^{2}$).
+For large displacements, the Green-Lagrange strain tensor $\boldsymbol{E}$ depends on the unknown displacement field $\mathbf{u}$ and is defined as:
 $\boldsymbol{E}=\frac{1}{2}\left(\nabla\mathbf{u}+\nabla\mathbf{u}^{T}+\nabla\mathbf{u}^{T}\nabla\mathbf{u}\right).$ (1)
-In addition, this equation relates to the Cauchy-Green deformation tensor $\boldsymbol{C}$ by: $\boldsymbol{C}=2\boldsymbol{E}-\mathbf{I}$, where $\mathbf{I}$ is the identity tensor.
-For our nonlinear simulations, we use a two-parameter Mooney-Rivlin model, where the strain energy density can be written as functions of the Cauchy-Green deformation tensor invariants ($\boldsymbol{I_{i}(C)}$, with $i=1,2,3$) [10]:
-$\boldsymbol{W}=C_{01}\left(\boldsymbol{I_{1}(C)}\boldsymbol{I_{3}(C)}^{\frac{-1}{3}}-3\right)+C_{10}\left(\boldsymbol{I_{1}(C)}\boldsymbol{I_{3}(C)}^{\frac{-2}{3}}-3\right),$ (8)
+This equation relates to the Cauchy-Green deformation tensor $\boldsymbol{C}$ by: $\boldsymbol{C}=2\boldsymbol{E}-\mathbf{I}$, where $\mathbf{I}$ is the identity tensor.
+For our nonlinear simulations, we use a two-parameter Mooney-Rivlin model [31], where the strain energy density can be written as functions of the Cauchy-Green deformation tensor invariants ($\boldsymbol{I_{i}(C)}$, with $i=1,2,3$) [13]:
+$\boldsymbol{W}=C_{01}\left(I_{1}(\mathbf{C})\cdot I_{3}(\mathbf{C})^{-1/3}-3\right)\\ +C_{10}\left(I_{1}(\mathbf{C})\cdot I_{3}(\mathbf{C})^{-2/3}-3\right),$ (2)
 where $C_{10}=1.62$ kPa and $C_{01}=1.97$ kPa are constitutive constants derived from ex-vivo porcine liver experiments [4].
@@
 Given a mesh partitioned into regions $\Omega_{i}$ informed by cholecystectomy surgery, nodes $\mathbf{x}_{s}$ are sampled from a mixture of Generalized von Mises-Fisher (vMF) distributions (i.e., radial densities on the surface mesh) centered in these regions:
-$p(\mathbf{x}_{s})=\sum_{i}g(\Omega_{i})f_{\text{vMF}}(\mathbf{x}_{s}\mid\boldsymbol{\mu}_{i},\kappa_{i}),\vspace{-.1in}$ (10)
-where $g(\Omega_{i})$ is the prior-informed probability of selecting $\Omega_{i}$, $\boldsymbol{\mu}_{i}$ indicates the center of $\Omega_{i}$, and $\kappa_{i}$ is the spread of the vMF, which is a function of the size of $\Omega_{i}$.
-The signals $u_{s}$ are independently sampled from a uniform distribution in a cuboid $\mathcal{C}$ defined by: $x\in(x_{1},x_{2}),y\in(y_{1},y_{2}),z\in(z_{1},z_{2})$, with volume: $V=(x_{2}-x_{1})(y_{2}-y_{1})(z_{2}-z_{1}).$ This approach ensures that node sampling is spatially structured according to surgical priors while allowing displacements to be isotropic in a bounded range.
+$p(\mathbf{x}_{s})=\sum_{i}g(\Omega_{i})f_{\text{vMF}}(\mathbf{x}_{s}\mid\boldsymbol{\mu}_{i},\kappa_{i}),$ (4)
+where $g(\Omega_{i})$ is the prior-informed probability of selecting $\Omega_{i}$, $\boldsymbol{\mu}_{i}$ indicates the center of $\Omega_{i}$, and $\kappa_{i}$ is the spread of the vMF, which depends on the size of $\Omega_{i}$.
+The displacements $\mathbf{u}_{s}\in\mathbb{R}^{3}$ are independently sampled from a uniform distribution over a cuboid domain $\mathcal{C}\subset\mathbb{R}^{3}$, defined by:
+$x\in(x_{1},x_{2}),\quad y\in(y_{1},y_{2}),\quad z\in(z_{1},z_{2}),$
+with total volume $V=(x_{2}-x_{1})(y_{2}-y_{1})(z_{2}-z_{1})$.
+This ensures that node sampling is spatially structured according to surgical priors, while displacements remain isotropic within a bounded range.
 For single grasping, ten regions $\Omega_{i}$ around the liver edge were selected, each with 250 samples.
 For multiple graspers, six pairwise combinations of five regions $\Omega_{i}$ were considered, with 30 samples per region.
-The displacement ranges are $x\in(x_{1},x_{2})=(-90,50)$ mm, $y\in(y_{1},y_{2})=(-90,50)$ mm, and $z\in(z_{1},z_{2})=(-60,90)$ mm, corresponding to the cube domain.
+The displacement bounds are $x\in(-90,50)$ $m$, $y\in(-90,50)$ mm, and $z\in(-60,90)$ $m$.
 2.0.2 Distribution for Multiple Graspers.
-For multiple interactions (e.g., two graspers), we define the joint distribution as $p(\mathbf{x}_{s}^{1},\mathbf{x}_{s}^{2})=p(\mathbf{x}_{s}^{1})p(\mathbf{x}_{s}^{2}\mid\mathbf{x}_{s}^{1})$, where $p(\mathbf{x}_{s}^{1})$ follows Eq.
+For multiple interactions (e.g., two graspers), we define the joint distribution as $p(\mathbf{x}_{1},\mathbf{x}_{2})=p(\mathbf{x}_{1})\,p(\mathbf{x}_{2}\mid\mathbf{x}_{1}),$ where $p(\mathbf{x}_{1})$ follows Eq.
 (4).
-To ensure that $\mathbf{x}_{s}^{2}$ is not sampled from the same region $\Omega_{k}$ as $\mathbf{x}_{s}^{1}$, we modify the region probabilities $g(\Omega_{i})$ by setting $g(\Omega_{k})=0$ and adjusting $g(\Omega_{i\neq k})$ based on surgical priors, and then sampling $\mathbf{x}_{s}^{2}$ from $p(\mathbf{x}_{s}^{2}\mid\mathbf{x}_{s}^{1})$.
+To ensure that $\mathbf{x}_{2}$ is not sampled from the same region $\Omega_{k}$ as $\mathbf{x}_{1}$, we modify the region probabilities $g(\Omega_{i})$ by setting $g(\Omega_{k})=0$, followed by renormalizing the remaining weights $g(\Omega_{i})$ for $i\neq k$ based on surgical priors.
+The conditional distribution $p(\mathbf{x}_{2}\mid\mathbf{x}_{1})$ is then defined using these updated region weights.
 2.0.3 Regularization Distribution, $q$.
-For Kelvinlet regularization, interaction points are sampled uniformly from the surface nodes.
-Initial displacements are then drawn uniformly from a cone centered around the surface normal at $\mathbf{x}_{s}$, with a maximum opening angle of $\alpha_{\max}$.
-4 Texture Rendering for Realistic Simulations
-Our visualization pipeline combines real-time tissue deformation and high-fidelity rendering to create a realistic simulation (see Figure 1 and the supplementary videos).
-In a preprocessing stage, surface meshes are extracted from the segmented high-resolution volumetric mesh, assigned texture coordinates, and enriched with procedurally generated textures using Worley noise [26].
-Interpolation weights are computed to map low-resolution deformations to the high-resolution mesh [16], optimized using a Bounding Volume Hierarchy (BVH) [5], and UV coordinates are transferred to the volumetric mesh.
-At runtime, the system dynamically updates surface meshes to accommodate added or removed tetrahedral elements, interpolates high-resolution vertex positions based on low-resolution deformations, smooths exposed internal structures, and recalculates surface normals.
-Rendering employs pre-integrated diffuse shading [19], triplanar texture projection for inner tissue surfaces [7], and post-processing effects within Unity to achieve realistic tissue appearance.
-The pipeline achieves near-real-time performance while preserving visual fidelity during dynamic deformations.
+For Kelvinlet regularization, interaction points $\mathbf{x}_{s}$ are sampled uniformly from the surface nodes.
+The corresponding displacements $\mathbf{u}_{s}\in\mathbb{R}^{3}$ are then sampled uniformly from a cone centered around the surface normal at $\mathbf{x}_{s}$, with an opening angle of $\alpha_{\max}$.
+Having detailed the data generation process, we now turn to our proposed neural augmented Kelvinlet solution.
+3 Neural-Augmented Kelvinlet
+In this section, we present the Kelvinlets prior and our residual learning and regularization frameworks.
+3.1 Navier-Cauchy Eqs and Green’s Function Solution
+Figure 2: Overview of the two proposed training frameworks.
+Method 1 (Left): Residual learning predicts only the deviation from the Kelvinlet solution, reducing learning complexity by leveraging the physics-based analytical solution as an initialization.
+Method 2 (Right): Kelvinlet-based regularization combines FEM-supervised learning with a secondary loss enforcing consistency with Kelvinlet priors.
+The heatmap depicts the displacement norms.
+The displacement field $\mathbf{u}:\mathbb{R}^{3}\to\mathbb{R}^{3}$ of an isotropic, homogeneous elastic medium under an applied body force density $\mathbf{f}:\mathbb{R}^{3}\to\mathbb{R}^{3}$ satisfies the Navier-Cauchy equations:
+$\mu\nabla^{2}\mathbf{u}+\frac{\mu}{1-2\nu}\nabla(\nabla\cdot\mathbf{u})+\mathbf{f}=0,$ (5)
+where $\mu>0$ is the shear modulus, and $\nu\in(0,0.5)$ is the Poisson’s ratio, ensuring a physically consistent material response.
+Here, $\nabla$ denotes the gradient operator, $\nabla\cdot\mathbf{u}$ is the divergence of the displacement field, and $\nabla^{2}\mathbf{u}=\nabla\cdot(\nabla\mathbf{u})$ is the vector Laplacian applied to $\mathbf{u}$.
+For a concentrated force $\mathbf{f}\delta(\mathbf{x}-\mathbf{x}_{s})$ applied at a source point $\mathbf{x}_{s}\in\mathbb{R}^{3}$, the Green’s function solution, known as the Kelvinlet, provides an analytical displacement field $\mathbf{u}:\mathbb{R}^{3}\setminus\{\mathbf{x}_{s}\}\to\mathbb{R}^{3}$ following:
+$\begin{split}\mathbf{u}(\mathbf{x})=\frac{1}{4\pi\mu}\Bigg(\frac{(1-\nu)I}{r}+\nu\frac{(\mathbf{x}-\mathbf{x}_{s})\otimes(\mathbf{x}-\mathbf{x}_{s})}{r^{3}}\Bigg)\mathbf{f}\end{split}$ (6)
+where $r=\|\mathbf{x}-\mathbf{x}_{s}\|_{2}$ denotes the Euclidean distance from the source, and $\otimes$ denotes the outer product operator.
+The classical Kelvinlet in Eq.
+(6) exhibits singularities at $\mathbf{x}_{s}$.
+To prevent numerical instabilities, we substitute $r$ with $r_{\varepsilon}=\sqrt{\|\mathbf{x}-\mathbf{x}_{s}\|^{2}+\varepsilon^{2}}$ for $\epsilon>0$, where the spatial extent of force distribution is controlled by $\varepsilon$, ensuring the solution is well-conditioned.
+Finally, a re-scaled solution with respect to the local displacement boundary condition $\mathbf{u}_{s}=\mathbf{u}(\mathbf{x}_{s})$ removes the dependence on both the force $\mathbf{f}$ and the modulus $\mu$, yielding the regularized Kelvinlet formulation used in this paper,
+$\mathbf{u}_{\varepsilon}(\mathbf{x})=\frac{\varepsilon}{(5-6\nu)}\Bigg(\frac{(3-4\nu)I}{r_{\varepsilon}}+\frac{(\mathbf{x}-\mathbf{x}_{s})\otimes(\mathbf{x}-\mathbf{x}_{s})}{r_{\varepsilon}^{3}}\Bigg)\mathbf{u}_{s}.$ (7)
+For multiple interaction points, we note that at each interaction point we only observe the superposition of contributions from all interaction points, without knowing the exact contribution of each individual point to this superposition.
+Consequently, unlike the single-grasper setting, $\mathbf{u}_{s}$ in Eq.
+(7) is not directly available for each interaction point and must be estimated.
+To formalize this, let $\mathbf{x}_{i}$ denote the location of the $i$th interaction, $\mathbf{u}_{i}$ the corresponding total displacement at $\mathbf{x}_{i}$, and $\mathbf{k}_{i}$ the unknown contribution of the $i$th interaction to this displacement.
+Assuming there are $K$ interaction points, the total displacement can be expressed as:
+$\mathbf{u}_{\epsilon}(\mathbf{x})=\sum_{i=1}^{K}\underbrace{\frac{\varepsilon}{(5-6\nu)}\left(\frac{(3-4\nu)I}{r_{\varepsilon}}+\frac{(\mathbf{x}-\mathbf{x}_{i})\otimes(\mathbf{x}-\mathbf{x}_{i})}{r_{\varepsilon}^{3}}\right)}_{\Gamma_{i}(\mathbf{x})}\mathbf{k_{i}},$ (8)
+where $\Gamma_{i}(\mathbf{x})\in\mathbb{R}^{3\times 3}$. Note that this formulation enables the total displacement function to incorporate cross communication between interaction points.
+To estimate the $\mathbf{k}_{i}$ values, we solve the following optimization problem, which enforces the boundary conditions $\mathbf{u}_{\varepsilon}(\mathbf{x}_{i})=\mathbf{u}_{i}$ for all $i$:
+$\operatorname*{argmin}_{\mathbf{k}}\sum_{i=1}^{K}\left(\|\sum_{j=1}^{K}\Gamma_{j}(\mathbf{x}_{i})\mathbf{k}_{j}-\mathbf{u}_{i}\|^{2})+\lambda\|\mathbf{k}_{i}\|^{2}\right).$ (9)
+This formulation ensures that the individual Kelvinlet contributions collectively satisfy the prescribed displacements, while the regularization term $\lambda=0.001$ promotes a minimum-energy solution.
+The optimization in Eq.
+(9) is essentially a regularized least squares problem, which admits a fast analytic solution:
+$\mathbf{k}^{*}=\left(\mathbf{\Gamma}^{\top}\mathbf{\Gamma}+\lambda\mathbf{I}\right)^{-1}\mathbf{\Gamma}^{\top}\mathbf{u},$ (10)
+where $\mathbf{I}\in\mathbb{R}^{3K\times 3K}$ is the identity matrix, and $\mathbf{\Gamma}\in\mathbb{R}^{3K\times 3K}$, $\mathbf{k}\in\mathbb{R}^{3K}$, and $\mathbf{u}\in\mathbb{R}^{3K}$ are defined as:
+$\mathbf{\Gamma}=\begin{bmatrix}\Gamma_{1}(\mathbf{x}_{1})&\cdots&\Gamma_{K}(\mathbf{x}_{1})\\ \vdots&\ddots&\vdots\\ \Gamma_{1}(\mathbf{x}_{K})&\cdots&\Gamma_{K}(\mathbf{x}_{K})\end{bmatrix},\mathbf{k}=\begin{bmatrix}\mathbf{k}_{1}\\ \vdots\\ \mathbf{k}_{K}\end{bmatrix},\mathbf{u}=\begin{bmatrix}\mathbf{u}_{1}\\ \vdots\\ \mathbf{u}_{K}\end{bmatrix}$.
+Note that while the closed form solution requires the inversion of a $3K\times 3K$ matrix, the number of interaction points, $K$ is often small, leading to a very fast solution for Kelvinlet.
+The total Kelvinlet deformation model, including the analytic solution to the optimization, can be performed in less than 1 ms for two simultaneously engaged graspers (i.e., $K\approx 20$) using native GPU acceleration in PyTorch on an NVIDIA A6000 GPU and on a mesh with $\sim 10\text{k}$ nodes.
+Having established a fast physics-based solution, we now focus on integrating Kelvinlets into our models.
+3.2 Method 1: Kelvinlet-Based Residual Learning
+For an initial displacement $\mathbf{u}_{s}$ located at $\mathbf{x}_{s}$, let $\mathbf{u}_{\text{true}}(\cdot;\mathbf{x}_{s},\mathbf{u}_{s})$ denote the ground-truth displacement field.
+Instead of directly regressing $\mathbf{u}_{\text{true}}$, we introduce a residual learning framework where a neural network estimates the deviation from the Kelvinlet prior, which serves as an initial approximation.
+Specifically, we define the residual displacement as $\mathbf{r}(\mathbf{x};\mathbf{x}_{s},\mathbf{u}_{s})=\mathbf{u}_{\text{true}}(\mathbf{x};\mathbf{x}_{s},\mathbf{u}_{s})-\mathbf{u}_{\varepsilon}(\mathbf{x};\mathbf{x}_{s},\mathbf{u}_{s}),$ where $\mathbf{u}_{\varepsilon}$ represents the Kelvinlet solution.
+The neural network, parameterized by $\boldsymbol{\theta}\in\mathbb{R}^{d}$, learns to approximate this residual via $\hat{\mathbf{r}}(\mathbf{x};\mathbf{x}_{s},\mathbf{u}_{s},\boldsymbol{\theta}).$ The corrected displacement field is then given by $\tilde{\mathbf{u}}(\mathbf{x};\mathbf{x}_{s},\mathbf{u}_{s},\boldsymbol{\theta})=\mathbf{u}_{\varepsilon}(\mathbf{x};\mathbf{x}_{s},\mathbf{u}_{s})+\hat{\mathbf{r}}(\mathbf{x};\mathbf{x}_{s},\mathbf{u}_{s},\boldsymbol{\theta}).$ Let $\|\mathbf{u}\|_{L_{2}(\Omega)}^{2}=\int_{\Omega}\|\mathbf{u}(\mathbf{x})\|^{2}d\mathbf{x}$, then to train the network, we minimize the loss:
+$\mathcal{L}_{\text{residual}}(\boldsymbol{\theta})=\mathbb{E}_{(\mathbf{x}_{s},\mathbf{u}_{s})\sim p}\Big[\|\mathbf{u}_{\text{true}}(\cdot;\mathbf{x}_{s},\mathbf{u}_{s})\\ -\tilde{\mathbf{u}}(\cdot;\mathbf{x}_{s},\mathbf{u}_{s},\boldsymbol{\theta})\|^{2}_{L_{2}(\Omega)}\Big],$ (11)
+where $p$ represents the distribution of all possible pointwise interactions with the tissue, and $\Omega$ denotes the tissue domain.
+By focusing on learning only the correction beyond the Kelvinlet solution, this approach leverages the physics priors directly into its formulation.
+3.3 Method 2: Kelvinlets as Regularization
+Alternatively, a network can be trained to directly regress $\mathbf{u}_{\text{true}}$, producing an estimate $\hat{\mathbf{u}}$.
+To enforce physically consistent predicted deformations by the network, we leverage Kelvinlets as a regularization prior and write:
+$\mathcal{L}(\boldsymbol{\theta})=\mathbb{E}_{(\mathbf{x}_{s},\mathbf{u}_{s})\sim p}\left[\left\|\mathbf{u}_{\text{true}}-\hat{\mathbf{u}}\right\|^{2}_{L_{2}(\Omega)}\right]\\ +\lambda_{\text{reg}}\,\mathbb{E}_{(\mathbf{x}^{\prime}_{s},\mathbf{u}^{\prime}_{s})\sim q}\left[\left\|\mathbf{u}_{\varepsilon}-\hat{\mathbf{u}}\right\|^{2}_{L_{2}(\Omega)}\right]$ (12)
+where $\lambda_{\text{reg}}>0$ is the regularization coefficient.
+For clarity and brevity, we omit the dependencies $(\cdot;\mathbf{x}_{s},\mathbf{u}_{s})$ and $(\cdot;\mathbf{x}_{s},\mathbf{u}_{s},\boldsymbol{\theta})$ in the notation.
+Note that we explicitly separate $(\mathbf{x}_{s},\mathbf{u}_{s})$ from $(\mathbf{x}_{s}^{\prime},\mathbf{u}^{\prime}_{s})$ because, in practice, the two distributions $p$ and $q$ could differ, and also while Monte Carlo integration is used for both terms, the first term requires FEM simulations to obtain $\mathbf{u}_{\text{true}}$ (which is computationally expensive), whereas the second term only involves Kelvinlet evaluations (much faster).
+We will describe the distributions $p$ and $q$ used in our experiments in the following section.
+Figure 2 illustrates the proposed residual- and regularization-based learning methods, along with the displacement norm visualizations for all methods on a representative sample.
 4 Experiments
 4.0.1 Data.
+Figure 3: Qualitative comparison of deformation predictions for the best (PointNet) and worst (GraphGPS) models in linear regimes.
 Our dataset consists of large-scale FEM simulations of soft tissue deformation.
@@
 4.0.2 Neural Architectures.
-We evaluate Kelvinlet-based priors using four neural architectures: U-Net [15], SAB [13] (set attention), GraphGPS [23] (a Graph Neural Network and Transformer hybrid), and PointNet [21] (a universal function approximator for sets).
-We modify PointNet and GraphGPS by reducing layers, U-Net by replacing convolutions with shared MLPs to form a DeepSet, and SAB by using two layers with an MLP regressor head.
-Each model is trained separately on two datasets: one with individual graspers and another with multiple graspers interacting.
-For each dataset, models are trained under three settings: (1) a baseline without physics priors (i.e., evaluating solely the network), (2) residual learning predicting deviations from Kelvinlets, and (3) physics-informed regularization enforcing Kelvinlet constraints.
+We first evaluate several deformation-specific baselines designed for soft tissue modeling.
+PhysGNN [33] is a mesh-based GNN trained on FEM simulations; we enhance it with a global attention layer after each graph convolution to mitigate the well-known oversmoothing associated with GNNs [35].
+PEGNN [32] models contact dynamics via cross-graph attention; we retain only its soft-tissue encoder to isolate local geometric features.
+For SeeSaw [9], we reimplement the method and adapt it to a point-cloud-based simulation setup.
+From the diffeomorphic mapping approach [12], we extract only the deformation network to enable direct supervision.
+We also include a Transformer-based KAN [16], limited to single-step prediction for fair comparison.
+To assess architecture-agnostic effects of Kelvinlet priors, we additionally evaluate U-Net [22], SAB [19], GraphGPS [29], and PointNet [28].
+U-Net is converted into a DeepSet by replacing convolutions with shared MLPs.
+SAB uses two self-attention layers and an MLP head.
+GraphGPS and PointNet are simplified by reducing their depth to prevent overfitting on high-resolution meshes.
+Figure 4: Log-scaled inference time vs.
+accuracy (DCM%) for various models in linear and nonlinear simulations.
+Transparent markers denote base models, while solid markers indicate Kelvinlet-regularized versions.
+Black-edged markers represent multi-grasper cases, while edge-free markers denote single-grasper cases.
+FEM and Kelvinlet baselines are included for reference.
 4.0.3 Hyperparameters.
@@
 4.1 Results and Analysis
-Figure 2: Qualitative comparison of deformation predictions for the best (PointNet) and worst (GraphGPS) models in linear (top) and nonlinear (bottom) regimes.
-To quantitatively assess model performance, we introduce the Deformation Capture Mean (DCM) metric, which measures how well a predicted deformation field approximates the ground-truth displacement field.
+Table 1: Deformation Capture Mean (DCM) (%) on individual graspers in linear simulations for modified baselines and our method, with and without Kelvinlet priors.
+Ours
+Model Base +Res. +Reg.
+PhysGNN[33] 90.12 (0.13) 91.77 (0.11) 92.65 (0.10)
+SeeSaw[9] 89.24 (0.15) 91.02 (0.10) 91.83 (0.08)
+Physics-Encoded GNN[32] 88.55 (0.14) 90.44 (0.09) 91.58 (0.12)
+Transformer-KAN[16] 90.89 (0.12) 92.01 (0.08) 93.22 (0.07)
+DiffoMorphic Deformation[12] 91.14 (0.10) 92.88 (0.11) 94.33 (0.09)
+PointNet[28] 94.23 (0.11) 94.66 (0.09) 96.54 (0.09)
+U-Net[22] 92.11 (0.12) 93.56 (0.10) 94.04 (0.11)
+GraphGPS[29] 88.60 (0.21) 89.45 (0.15) 89.61 (0.23)
+SAB[19] 93.42 (0.11) 94.60 (0.04) 96.03 (0.04)
+Kelvinlet 65.24 (0) - -
+Average Gain +1.23 +2.39
+Minimum/Maximum Gain +0.43/1.89 +1.01/3.19
+Table 2: Inference time (ms) on an NVIDIA RTX A6000 GPU for a single grasper with individual models in linear simulations, with and without Kelvinlet priors.
+Inference time (ms) Ours
+Model Base +Res. +Reg.
+PhysGNN[33] 1.71 (0.11) 1.84 (0.13) 1.71 (0.11)
+SeeSaw[9] 1.56 (0.09) 1.69 (0.08) 1.56 (0.09)
+Physics-Encoded GNN[32] 1.65 (0.08) 1.79 (0.08) 1.65 (0.08)
+Transformer-KAN[16] 2.22 (0.17) 2.32 (0.16) 2.22 (0.17)
+DiffoMorphic Deformation[12] 2.16 (0.12) 2.35 (0.11) 2.16 (0.12)
+PointNet[28] 1.16 (0.10) 1.23 (0.13) 1.16 (0.10)
+U-Net[22] 1.55 (0.08) 1.64 (0.07) 1.55 (0.08)
+GraphGPS[29] 2.60 (0.15) 2.78 (0.15) 2.60 (0.15)
+SAB[19] 2.30 (0.02) 2.44 (0.21) 2.30 (0.20)
+Kelvinlet 0.15 (0.052)
+Table 3: Performance of various models on linear simulations with individual graspers, combined graspers, and multitask learning.
+DCM (%) is reported for single-task and multi-task settings with standard deviations.
+Model Indiv. Graspers (DCM) Comb. Graspers (DCM) Multitask (DCM)
+Base +Res. +Reg. Base +Res. +Reg. Base +Res. +Reg.
+U -Net[22] 92.11 (0.12) 93.56 (0.10) 94.04 (0.11) 88.23 (0.08) 89.67 (0.09) 90.01 (0.08) 91.20 (0.09) 92.33 (0.10) 92.88 (0.08)
+GraphGPS[29] 88.60 (0.21) 89.45 (0.15) 89.61 (0.23) 87.33 (0.06) 87.92 (0.02) 88.01 (0.04) 88.35 (0.13) 88.91 (0.10) 89.45 (0.09)
+PointNet[28] 94.23 (0.11) 94.66 (0.09) 96.54 (0.09) 89.79 (0.07) 90.11 (0.07) 91.22 (0.07) 94.00 (0.08) 94.12 (0.07) 95.01 (0.06)
+SAB[19] 93.42 (0.11) 94.60 (0.04) 96.03 (0.04) 88.05 (0.09) 89.91 (0.11) 91.02 (0.04) 92.54 (0.06) 93.66 (0.07) 94.42 (0.06)
+Kelvinlet 65.24 (0) 64.33 (0) 64.82 (0)
+Table 4: Performance of various models on nonlinear simulations with individual graspers, combined graspers, and multitask learning.
+DCM (%) is reported for single-task and multi-task settings with standard deviations.
+Model Indiv. Graspers (DCM) Comb. Graspers (DCM) Multitask (DCM)
+Base +Res. +Reg. Base +Res. +Reg. Base +Res. +Reg.
+U -Net[22] 80.56 (0.10) 84.22 (0.09) 84.88 (0.11) 76.91 (0.05) 79.83 (0.08) 82.10 (0.21) 78.10 (0.12) 80.45 (0.09) 81.33 (0.14)
+GraphGPS[29] 77.52 (0.17) 80.63 (0.13) 82.11 (0.09) 73.72 (0.15) 77.92 (0.10) 80.66 (0.16) 76.24 (0.18) 79.31 (0.15) 80.42 (0.10)
+PointNet[28] 81.25 (0.22) 84.65 (0.19) 85.05 (0.15) 79.56 (0.18) 82.66 (0.16) 83.33 (0.11) 83.25 (0.17) 84.75 (0.13) 85.21 (0.09)
+SAB[19] 80.75 (0.19) 84.44 (0.07) 84.75 (0.10) 78.88 (0.11) 81.50 (0.15) 82.83 (0.16) 81.22 (0.10) 83.01 (0.08) 84.00 (0.07)
+Kelvinlet 58.88 (0) 51.67 (0) 55.56 (0)
+To quantitatively assess model performance, we utilize the Deformation Capture Mean (DCM) metric, which measures how well a predicted deformation field approximates the ground-truth displacement field.
 The DCM is computed as:
 $\text{DCM}=100\times\left(1-\mathbb{E}_{(\mathbf{x}_{s},\mathbf{u}_{s})\sim p}\left[\frac{\left\|\mathbf{u}_{\text{true}}-\hat{\mathbf{u}}\right\|_{L_{2}(\Omega)}}{\|\mathbf{u}_{\text{true}}\|_{L_{2}(\Omega)}}\right]\right).$ (13)
-Tables 1 and 2 report DCMs across models and settings.
-Each experiment runs with three random seeds, reporting mean and standard deviation.
+Table 1 shows that integrating Kelvinlet-based priors improves deformation accuracy across all models.
+Residual learning yields an average gain of +1.23%, while Kelvinlet regularization achieves a stronger improvement of +2.39%, with the highest gain of +3.19% observed in DiffoMorphic Deformation.
+PointNet benefits the most in absolute performance, reaching 96.54% DCM.
+The analytical Kelvinlet alone performs poorly (65.24%), underscoring the value of learning residuals.
+As seen in Table 2, our methods add minimal runtime overhead.
+Regularized variants match the base inference time, while residual versions increase latency by only $\sim$0.1ms.
+Kelvinlet inference alone is extremely fast (0.15ms) but not accurate enough, justifying its role as a prior.
+Overall, our approach enhances accuracy without compromising efficiency.
+Tables 3 and 4 report DCMs across models and settings.
+Each experiment runs with three random seeds, reporting means and standard deviations.
 Models perform better on linear cases, as nonlinear deformations introduce gravity effects, non-proportional stress-strain relationships, and greater sensitivity to perturbations.
 Notably, Kelvinlet priors consistently improve DCMs, despite being derived for linear elasticity, showing their effectiveness even in nonlinear regimes.
-Table 1: DCM of models with/without Kelvinlet priors in linear simulations.
-Model Indiv. Graspers Comb. Graspers
-Base +Kelv. Res. +Kelv. Reg. Base +Kelv. Res. +Kelv. Reg.
-U-Net[15] 92.11 (0.12) 93.56 (0.10) 94.04 (0.11) 88.23 (0.08) 89.67 (0.09) 90.01 (0.08)
-GraphGPS[23] 88.60 (0.21) 89.45 (0.15) 89.61 (0.23) 87.33 (0.06) 87.92 (0.02) 88.01 (0.04)
-PointNet[21] 94.23 (0.11) 94.66 (0.09) 96.54 (0.09) 89.79 (0.07) 90.11 (0.07) 91.22 (0.07)
-SAB[13] 93.42 (0.11) 94.60 (0.04) 96.03 (0.04) 88.05 (0.09) 89.91 (011) 91.02 (0.04)
-Kelvinlet 65.24 (0) 64.33 (0)
-Table 2: DCM of models with/without Kelvinlet priors in nonlinear simulations.
-Model Indiv. Graspers Comb. Graspers
-Base +Kelv. Res. +Kelv. Reg. Base +Kelv. Res. +Kelv. Reg.
-U-Net[15] 80.56 (0.10) 84.22 (0.09) 84.88 (0.11) 76.91 (0.05) 79.83 (0.08) 82.10 (0.21)
-GraphGPS[23] 77.52 (0.17) 80.63 (0.13) 82.11 (0.09) 73.72 (0.15) 77.92 (0.10) 80.66 (0.16)
-PointNet[21] 81.25 (0.22) 84.65 (0.19) 85.05 (0.15) 79.56 (0.18) 82.66 (0.16) 83.33 (0.11)
-SAB[13] 80.75 (0.19) 84.44 (0.07) 84.75 (0.10) 78.88 (0.11) 81.50 (0.15) 82.83 (0.16)
-Kelvinlet 58.88 (0) 51.67 (0)
 Figure 4 illustrates the trade-off between inference time and accuracy for different models across linear and nonlinear simulations.
@@
 All computations were performed on a single NVIDIA A6000 GPU, ensuring a consistent hardware environment for runtime measurements.
-Figure 3: Log-scaled inference time vs.
-accuracy (DCM%) for various models in linear and nonlinear simulations.
-Transparent markers denote base models, while solid markers indicate Kelvinlet-regularized versions.
-Black-edged markers represent multi-grasper cases, while edge-free markers denote single-grasper cases.
-FEM and Kelvinlet baselines are included for reference.
 5 Conclusion
@@
 Our GPU-accelerated PyTorch implementation enables real-time inference, making it well-suited for surgical robotics and medical training, while a high-fidelity rendering pipeline ensures realistic soft tissue simulations in real time.
-Acknowledgment
-This work was partially supported by the NSF CAREER Award #2339898 and the Wellcome Leap ‘Surgery: Assess/Validate/Expand (SAVE)’ program.
+Acknowledgments
+This project is supported by the Wellcome Leap Surgery: Assess/Validate/Expand (SAVE) program.

### v3

@@
 These results establish Kelvinlet-augmented learning as a principled and computationally efficient paradigm for real-time, physics-aware soft-tissue simulation in surgical AI.
-Our code and data is available at: https://github.com/Anon92373/Neural-Kelvinlet.
+Code and datahttps://github.com/mint-vu/neural-augmented-kelvinlet-leap
+Introduction
 Soft tissue deformation modeling is essential in computational biomechanics for surgical simulation, image registration, and robotic-assisted interventions.
@@
 Traditional approaches rely on physics-based methods grounded in continuum mechanics, such as the Finite Element Method (FEM), Boundary Element Method (BEM), and Tensor Mass Model (TMM).
-While highly accurate and physically realistic, these methods are often computationally prohibitive for real-time applications [2, 34].
-To reduce computational cost, approximate methods such as mass-spring systems [23], mesh-free methods [3], and particle-based methods [21] are frequently used.
+While highly accurate and physically realistic, these methods are often computationally prohibitive for real-time applications (Allard, Courtecuisse, and Faure 2012; Zhang, Zhong, and Gu 2019).
+To reduce computational cost, approximate methods such as mass-spring systems (Nedel and Thalmann 1998), mesh-free methods (Belytschko et al.
+1996), and particle-based methods (Liu et al.
+2015) are frequently used.
 However, these approximations often sacrifice realism, creating a significant simulation-to-reality gap that limits their practicality in high-fidelity applications.
 Recent advances in deep learning and neural operators have enabled fully data-driven methods that approximate soft tissue deformations with a single forward pass of a learned model.
-Convolutional neural networks (CNNs), graph neural networks (GNNs), and Transformers [26] have demonstrated the ability to capture complex deformation patterns from simulated or real-world data, enabling real-time tissue simulation.
+Convolutional neural networks (CNNs), graph neural networks (GNNs), and Transformers (Wu, Kazanzides, and Unberath 2020; Wu et al.
+2021; Pfaff et al.
+2021) have demonstrated the ability to capture complex deformation patterns from simulated or real-world data, enabling real-time tissue simulation.
 Despite these advances, fully data-driven approaches face two fundamental limitations: (a) limited generalization to unseen force distributions and boundary conditions, which results in unreliable extrapolation, and (b) a tendency to produce physically inconsistent deformations, particularly when trained on datasets of limited size or diversity.
 These shortcomings highlight the need for hybrid approaches that combine the expressiveness of neural models with the robustness of physics-based priors.
-To address these challenges, considerable efforts have focused on integrating physics priors into data-driven machine learning models [17].
-Broadly, these methods make learning algorithms physics-informed by introducing: (a) observational biases, such as data augmentations that embed physical constraints; (b) inductive biases in the model architecture, for example, using group-invariant or equivariant neural networks [5]; and (c) learning biases through the design of loss functions that explicitly enforce physical laws.
+To address these challenges, considerable efforts have focused on integrating physics priors into data-driven machine learning models (Karniadakis et al.
+2021).
+Broadly, these methods make learning algorithms physics-informed by introducing: (a) observational biases, such as data augmentations that embed physical constraints; (b) inductive biases in the model architecture, for example, using group-invariant or equivariant neural networks (Bronstein et al.
+2017); and (c) learning biases through the design of loss functions that explicitly enforce physical laws.
 Despite this progress, achieving the necessary low latency for real-time simulation while maintaining a high-performing physics-informed model with appropriately chosen biases remains a challenging and largely open problem.
@@
 This is analogous to determining how the entire soft tissue will deform when graspers engage multiple surface locations and are moved to specified target positions.
-While several prior studies have explored similar setup and demonstrated that neural networks can approximate the solutions produced by high-fidelity physics-based solvers such as FEM [27], these approaches often lack strong physical priors and can suffer from reduced generalization and physical inconsistency.
+While several prior studies have explored similar setup and demonstrated that neural networks can approximate the solutions produced by high-fidelity physics-based solvers such as FEM (Pfeiffer et al.
+2019), these approaches often lack strong physical priors and can suffer from reduced generalization and physical inconsistency.
 To overcome these limitations, we propose a physics-informed neural framework that integrates analytical elastic priors to enhance both accuracy and physical plausibility.
-We introduce a physics-informed neural framework that leverages Kelvinlets [8], i.e., the closed-form Green’s function solutions to the Navier–Cauchy equations, as an analytically tractable and computationally efficient prior.
+We introduce a physics-informed neural framework that leverages Kelvinlets (De Goes and James 2017), i.e., the closed-form Green’s function solutions to the Navier–Cauchy equations, as an analytically tractable and computationally efficient prior.
 Kelvinlets provide a fast, low-order approximation of elastic deformation that captures the fundamental physics of linear elasticity.
@@
     We introduce two hybrid approaches that integrate Kelvinlets into data-driven soft tissue deformation modeling and conduct extensive experiments to demonstrate their superiority over purely physics-informed and data-driven baselines across diverse neural architectures.
-1 Related Work
+Related Work
 Classical approaches based on the Finite Element Method (FEM) provide high-fidelity simulations of soft tissue deformation but are often computationally prohibitive for real-time surgical applications.
@@
 For example, Lau et al.
-[20] employ a reduced-order linear FEM with modal warping to achieve sub-15 ms inference for liver deformation, enabling intraoperative use.
-Complementarily, Hu and Desai [14] characterize nonlinear liver material behavior through indentation experiments and validate Local Effective Moduli (LEM) estimates using hyperelastic FEM models, laying foundational work for realistic soft-tissue mechanics under large strains.
+(Liu et al.
+2025) employ a reduced-order linear FEM with modal warping to achieve sub-15 ms inference for liver deformation, enabling intraoperative use.
+Complementarily, Hu and Desai (Hu and Desai 2004) characterize nonlinear liver material behavior through indentation experiments and validate Local Effective Moduli (LEM) estimates using hyperelastic FEM models, laying foundational work for realistic soft-tissue mechanics under large strains.
 More recently, physics-informed neural networks (PINNs) have emerged as a data-efficient alternative that integrates governing equations such as elasticity or dynamics directly into the learning objective.
 Hu et al.
-[15] introduce a PINN that encodes static linear elasticity to predict 3D stress and strain fields of liver tissue under surgical tool interaction, achieving millisecond-scale inference while retaining physical plausibility.
+(Hu et al.
+2025) introduce a PINN that encodes static linear elasticity to predict 3D stress and strain fields of liver tissue under surgical tool interaction, achieving millisecond-scale inference while retaining physical plausibility.
 Extending this paradigm, Nguyen-Le et al.
-[24] propose a PINODE framework that augments neural dynamics with mass-spring priors, enabling extrapolation from sparse motion data.
+(Nguyen-Le, Ballit, and Dao 2025) propose a PINODE framework that augments neural dynamics with mass-spring priors, enabling extrapolation from sparse motion data.
 These methods improve interpretability and generalization but often require complex training and remain constrained by solver latency or limited scalability.
@@
 Ke et al.
-[6] train a feedforward network to predict liver surface displacements conditioned on retractor pose, achieving sub-millisecond inference.
+(Camara et al.
+2016) train a feedforward network to predict liver surface displacements conditioned on retractor pose, achieving sub-millisecond inference.
 While fast and simple, such models lack physical constraints, which can reduce reliability and robustness in unseen conditions.
-On the learning-based front, mesh-based models such as PhysGNN leverage FEM supervision to train graph neural networks for real-time deformation prediction in neurosurgery [33].
-PEGNN [32] extends this with contact-aware hybrid graph reasoning, enabling accurate predictions across surgical and manipulation scenarios.
-Meanwhile, SeeSaw [9] employs stereo video and self-supervised learning to estimate sparse 3D deformations using a GNN encoder.
+On the learning-based front, mesh-based models such as PhysGNN leverage FEM supervision to train graph neural networks for real-time deformation prediction in neurosurgery (Salehi and Giannacopoulos 2022).
+PEGNN (Saleh et al.
+2024) extends this with contact-aware hybrid graph reasoning, enabling accurate predictions across surgical and manipulation scenarios.
+Meanwhile, SeeSaw (Docea et al.
+2024) employs stereo video and self-supervised learning to estimate sparse 3D deformations using a GNN encoder.
 Zhao et al.
-[12] introduce a diffeomorphic mapping framework that enforces cycle consistency, capturing invertible tissue motion.
+(Gong et al.
+2024) introduce a diffeomorphic mapping framework that enforces cycle consistency, capturing invertible tissue motion.
 These approaches show strong empirical performance but may suffer from limited interpretability or reliance on dense supervision.
-Neural architectures such as U-Net [22], Set Attention Block (SAB) [19], and GraphGPS [29] offer modular and scalable backbones for modeling deformation fields across grid-based, set-based, and graph-structured data.
+Neural architectures such as U-Net (Mendizabal, Márquez-Neila, and Cotin 2020), Set Attention Block (SAB) (Lee et al.
+2019), and GraphGPS (Rampášek et al.
+2022) offer modular and scalable backbones for modeling deformation fields across grid-based, set-based, and graph-structured data.
 While effective in capturing spatial patterns and long-range dependencies, these models are generally agnostic to the underlying physics of soft tissue behavior and typically lack inductive biases grounded in biomechanics, such as force equilibrium, boundary conditions, or constitutive material priors.
@@
 This hybrid approach improves generalization, improves physical plausibility, and remains compatible with a range of backbone architectures, while preserving the low latency performance required for real-time surgical simulation and interactive applications.
-2 Large-Scale FEM Data
+Large-Scale FEM Data
 Figure 1: Overview of the large-scale FEM data generation pipeline.
@@
 Neural network–based solutions require training on large-scale finite element (FE) simulation datasets.
-In this work, we focus on laparoscopic cholecystectomy and construct an FE model from an abdominal CT image and its segmentation masks, converting them into a 3D linear tetrahedral mesh with 10,400 nodes and 44,045 tetrahedra using gmsh [11].
+In this work, we focus on laparoscopic cholecystectomy and construct an FE model from an abdominal CT image and its segmentation masks, converting them into a 3D linear tetrahedral mesh with 10,400 nodes and 44,045 tetrahedra using gmsh (Geuzaine and Remacle 2009).
 As part of our contributions, we release this large-scale dataset; the details of its creation are provided below.
@@
 These regions are identified by clustering nodes along the liver ridge, based on surgical video analysis and input from our surgical collaborator.
-FE simulations are performed using the open-source getFEM package [30], applying tool-imposed displacements at those predefined clustered regions.
+FE simulations are performed using the open-source getFEM package (Renard and Poulios 2020), applying tool-imposed displacements at those predefined clustered regions.
 This pipeline is illustrated in Figure 1.
@@
 This equation relates to the Cauchy-Green deformation tensor $\boldsymbol{C}$ by: $\boldsymbol{C}=2\boldsymbol{E}-\mathbf{I}$, where $\mathbf{I}$ is the identity tensor.
-For our nonlinear simulations, we use a two-parameter Mooney-Rivlin model [31], where the strain energy density can be written as functions of the Cauchy-Green deformation tensor invariants ($\boldsymbol{I_{i}(C)}$, with $i=1,2,3$) [13]:
+For our nonlinear simulations, we use a two-parameter Mooney-Rivlin model (Rivlin 1948), where the strain energy density can be written as functions of the Cauchy-Green deformation tensor invariants ($\boldsymbol{I_{i}(C)}$, with $i=1,2,3$) (Hackett 2018):
 $\boldsymbol{W}=C_{01}\left(I_{1}(\mathbf{C})\cdot I_{3}(\mathbf{C})^{-1/3}-3\right)\\ +C_{10}\left(I_{1}(\mathbf{C})\cdot I_{3}(\mathbf{C})^{-2/3}-3\right),$ (2)
-where $C_{10}=1.62$ kPa and $C_{01}=1.97$ kPa are constitutive constants derived from ex-vivo porcine liver experiments [4].
+where $C_{10}=1.62$ kPa and $C_{01}=1.97$ kPa are constitutive constants derived from ex-vivo porcine liver experiments (Benson, Pereira, and Miga 2025).
 For our linear simulations, we simulate an elastic response in the absence of gravitational load ($\boldsymbol{g}=0$).
@@
 $\boldsymbol{W}=\left(\frac{\mu\nu}{1-2\nu}\right)\mathrm{Tr}(\boldsymbol{E})^{2}+\mu\mathrm{Tr}(\boldsymbol{E}^{2}).$ (3)
-In these linear simulations, we assign a Poisson’s ratio of 0.45 for all organs [7].
-The shear modulus values are also set as follows: 0.03 kPa for the fascia [1], 0.34 kPa for the gallbladder [25], 0.69 kPa for the gallbladder wall and liver [18], and 1.07 kPa for the remaining organs [10].
-2.0.1 Training Distribution, $p$.
+In these linear simulations, we assign a Poisson’s ratio of 0.45 for all organs (Chen et al.
+1996).
+The shear modulus values are also set as follows: 0.03 kPa for the fascia (Akhmanova et al.
+2015), 0.34 kPa for the gallbladder (Nisansala et al.
+2016), 0.69 kPa for the gallbladder wall and liver (Kim et al.
+2013), and 1.07 kPa for the remaining organs (George et al.
+2018).
+Training Distribution, $p$.
 Given a mesh partitioned into regions $\Omega_{i}$ informed by cholecystectomy surgery, nodes $\mathbf{x}_{s}$ are sampled from a mixture of Generalized von Mises-Fisher (vMF) distributions (i.e., radial densities on the surface mesh) centered in these regions:
@@
 The displacement bounds are $x\in(-90,50)$ $m$, $y\in(-90,50)$ mm, and $z\in(-60,90)$ $m$.
-2.0.2 Distribution for Multiple Graspers.
+Distribution for Multiple Graspers.
 For multiple interactions (e.g., two graspers), we define the joint distribution as $p(\mathbf{x}_{1},\mathbf{x}_{2})=p(\mathbf{x}_{1})\,p(\mathbf{x}_{2}\mid\mathbf{x}_{1}),$ where $p(\mathbf{x}_{1})$ follows Eq.
@@
 The conditional distribution $p(\mathbf{x}_{2}\mid\mathbf{x}_{1})$ is then defined using these updated region weights.
-2.0.3 Regularization Distribution, $q$.
+Regularization Distribution, $q$.
 For Kelvinlet regularization, interaction points $\mathbf{x}_{s}$ are sampled uniformly from the surface nodes.
@@
 Having detailed the data generation process, we now turn to our proposed neural augmented Kelvinlet solution.
-3 Neural-Augmented Kelvinlet
+Neural-Augmented Kelvinlet
 In this section, we present the Kelvinlets prior and our residual learning and regularization frameworks.
-3.1 Navier-Cauchy Eqs and Green’s Function Solution
+Navier-Cauchy Eqs and Green’s Function Solution
 Figure 2: Overview of the two proposed training frameworks.
@@
 $\mathbf{u}_{\epsilon}(\mathbf{x})=\sum_{i=1}^{K}\underbrace{\frac{\varepsilon}{(5-6\nu)}\left(\frac{(3-4\nu)I}{r_{\varepsilon}}+\frac{(\mathbf{x}-\mathbf{x}_{i})\otimes(\mathbf{x}-\mathbf{x}_{i})}{r_{\varepsilon}^{3}}\right)}_{\Gamma_{i}(\mathbf{x})}\mathbf{k_{i}},$ (8)
-where $\Gamma_{i}(\mathbf{x})\in\mathbb{R}^{3\times 3}$. Note that this formulation enables the total displacement function to incorporate cross communication between interaction points.
+where $\Gamma_{i}(\mathbf{x})\in\mathbb{R}^{3\times 3}$.
+and $r_{i,\varepsilon}(\mathbf{x}):=\sqrt{\|\mathbf{x}-\mathbf{x}_{i}\|_{2}^{2}+\varepsilon^{2}}.$ Note that this formulation enables the total displacement function to incorporate cross communication between interaction points.
 To estimate the $\mathbf{k}_{i}$ values, we solve the following optimization problem, which enforces the boundary conditions $\mathbf{u}_{\varepsilon}(\mathbf{x}_{i})=\mathbf{u}_{i}$ for all $i$:
@@
 Having established a fast physics-based solution, we now focus on integrating Kelvinlets into our models.
-3.2 Method 1: Kelvinlet-Based Residual Learning
+Method 1: Kelvinlet-Based Residual Learning
 For an initial displacement $\mathbf{u}_{s}$ located at $\mathbf{x}_{s}$, let $\mathbf{u}_{\text{true}}(\cdot;\mathbf{x}_{s},\mathbf{u}_{s})$ denote the ground-truth displacement field.
@@
 By focusing on learning only the correction beyond the Kelvinlet solution, this approach leverages the physics priors directly into its formulation.
-3.3 Method 2: Kelvinlets as Regularization
+Method 2: Kelvinlets as Regularization
 Alternatively, a network can be trained to directly regress $\mathbf{u}_{\text{true}}$, producing an estimate $\hat{\mathbf{u}}$.
@@
 We will describe the distributions $p$ and $q$ used in our experiments in the following section.
-Figure 2 illustrates the proposed residual- and regularization-based learning methods, along with the displacement norm visualizations for all methods on a representative sample.
-4 Experiments
-4.0.1 Data.
+Figure  2 illustrates the proposed residual- and regularization-based learning methods, along with the displacement norm visualizations for all methods on a representative sample.
+Experiments
+Data.
 Figure 3: Qualitative comparison of deformation predictions for the best (PointNet) and worst (GraphGPS) models in linear regimes.
@@
 Of these, 5,000 correspond to single manipulations and 5,400 to multiple grasper manipulations.
-4.0.2 Neural Architectures.
-We first evaluate several deformation-specific baselines designed for soft tissue modeling.
-PhysGNN [33] is a mesh-based GNN trained on FEM simulations; we enhance it with a global attention layer after each graph convolution to mitigate the well-known oversmoothing associated with GNNs [35].
-PEGNN [32] models contact dynamics via cross-graph attention; we retain only its soft-tissue encoder to isolate local geometric features.
-For SeeSaw [9], we reimplement the method and adapt it to a point-cloud-based simulation setup.
-From the diffeomorphic mapping approach [12], we extract only the deformation network to enable direct supervision.
-We also include a Transformer-based KAN [16], limited to single-step prediction for fair comparison.
-To assess architecture-agnostic effects of Kelvinlet priors, we additionally evaluate U-Net [22], SAB [19], GraphGPS [29], and PointNet [28].
-U-Net is converted into a DeepSet by replacing convolutions with shared MLPs.
-SAB uses two self-attention layers and an MLP head.
-GraphGPS and PointNet are simplified by reducing their depth to prevent overfitting on high-resolution meshes.
+Neural Architectures.
+We evaluate two groups of models: deformation-specific FEM surrogates and generic set/graph architectures.
+For the former, PhysGNN (Salehi and Giannacopoulos 2022) is a mesh-based GNN that we augment with a global attention layer to counter oversmoothing (Zhao and Akoglu 2020); PEGNN (Saleh et al.
+2024) is used only through its soft-tissue encoder; SeeSaw (Docea et al.
+2024) is reimplemented for our point-cloud setting; from the diffeomorphic framework of (Gong et al.
+2024) we keep only the deformation network; and a Transformer-based KAN (Jiang et al.
+2025) is restricted to single-step prediction.
+To study architecture-agnostic effects of Kelvinlet priors, we also use U-Net (Mendizabal, Márquez-Neila, and Cotin 2020), SAB (Lee et al.
+2019), GraphGPS (Rampášek et al.
+2022), and PointNet (Qi et al.
+2016): U-Net is converted into a DeepSet-style model via shared pointwise MLPs, SAB uses two self-attention blocks with an MLP head, and GraphGPS and PointNet are shallow variants with reduced depth to avoid overfitting on high-resolution meshes.
 Figure 4: Log-scaled inference time vs.
@@
 FEM and Kelvinlet baselines are included for reference.
-4.0.3 Hyperparameters.
+Hyperparameters.
 We set shear modulus in Kelvinlet to $\mu=0.72$ kPa, Poisson’s ratio $\nu=0.45$, and regularization parameter $\epsilon=0.05$, to reflect soft tissue properties and ensure numerical stability.
 For linear and nonlinear simulations we use $\lambda_{\text{reg}}=1$ and $\lambda_{\text{reg}}=0.1$, respectively, to emphasize linearity of Kelvinlet.
-4.1 Results and Analysis
-Table 1: Deformation Capture Mean (DCM) (%) on individual graspers in linear simulations for modified baselines and our method, with and without Kelvinlet priors.
-Ours
-Model Base +Res. +Reg.
-PhysGNN[33] 90.12 (0.13) 91.77 (0.11) 92.65 (0.10)
-SeeSaw[9] 89.24 (0.15) 91.02 (0.10) 91.83 (0.08)
-Physics-Encoded GNN[32] 88.55 (0.14) 90.44 (0.09) 91.58 (0.12)
-Transformer-KAN[16] 90.89 (0.12) 92.01 (0.08) 93.22 (0.07)
-DiffoMorphic Deformation[12] 91.14 (0.10) 92.88 (0.11) 94.33 (0.09)
-PointNet[28] 94.23 (0.11) 94.66 (0.09) 96.54 (0.09)
-U-Net[22] 92.11 (0.12) 93.56 (0.10) 94.04 (0.11)
-GraphGPS[29] 88.60 (0.21) 89.45 (0.15) 89.61 (0.23)
-SAB[19] 93.42 (0.11) 94.60 (0.04) 96.03 (0.04)
-Kelvinlet 65.24 (0) - -
-Average Gain +1.23 +2.39
-Minimum/Maximum Gain +0.43/1.89 +1.01/3.19
-Table 2: Inference time (ms) on an NVIDIA RTX A6000 GPU for a single grasper with individual models in linear simulations, with and without Kelvinlet priors.
-Inference time (ms) Ours
-Model Base +Res. +Reg.
-PhysGNN[33] 1.71 (0.11) 1.84 (0.13) 1.71 (0.11)
-SeeSaw[9] 1.56 (0.09) 1.69 (0.08) 1.56 (0.09)
-Physics-Encoded GNN[32] 1.65 (0.08) 1.79 (0.08) 1.65 (0.08)
-Transformer-KAN[16] 2.22 (0.17) 2.32 (0.16) 2.22 (0.17)
-DiffoMorphic Deformation[12] 2.16 (0.12) 2.35 (0.11) 2.16 (0.12)
-PointNet[28] 1.16 (0.10) 1.23 (0.13) 1.16 (0.10)
-U-Net[22] 1.55 (0.08) 1.64 (0.07) 1.55 (0.08)
-GraphGPS[29] 2.60 (0.15) 2.78 (0.15) 2.60 (0.15)
-SAB[19] 2.30 (0.02) 2.44 (0.21) 2.30 (0.20)
+Results and Analysis
+Model Base Ours
++Res. +Reg.
+PhysGNN(Salehi and Giannacopoulos 2022) 1.71 (0.11) 1.84 (0.13) 1.71 (0.11)
+SeeSaw(Docea et al. 2024) 1.56 (0.09) 1.69 (0.08) 1.56 (0.09)
+Physics-Encoded GNN(Saleh et al. 2024) 1.65 (0.08) 1.79 (0.08) 1.65 (0.08)
+Transformer-KAN(Jiang et al. 2025) 2.22 (0.17) 2.32 (0.16) 2.22 (0.17)
+Diffeomorphic Deformation(Gong et al. 2024) 2.16 (0.12) 2.35 (0.11) 2.16 (0.12)
+PointNet(Qi et al. 2016) 1.16 (0.10) 1.23 (0.13) 1.16 (0.10)
+U-Net(Mendizabal, Márquez-Neila, and Cotin 2020) 1.55 (0.08) 1.64 (0.07) 1.55 (0.08)
+GraphGPS(Rampášek et al. 2022) 2.60 (0.15) 2.78 (0.15) 2.60 (0.15)
+SAB(Lee et al. 2019) 2.30 (0.02) 2.44 (0.21) 2.30 (0.20)
 Kelvinlet 0.15 (0.052)
+Table 1: Inference time (ms) on an NVIDIA RTX A6000 GPU for a single grasper in linear simulations, with and without Kelvinlet priors. Values are mean (std).
+Model Indiv. (DCM) Comb. (DCM) Multi (DCM)
+Base +Res. +Reg. Base +Res. +Reg. Base +Res. +Reg.
+PhysGNN(Salehi and Giannacopoulos 2022) 90.12(0.13) 91.77(0.11) 92.65(0.10) 86.38(0.13) 88.10(0.11) 88.66(0.10) 89.55(0.13) 90.96(0.11) 91.54(0.10)
+SeeSaw(Docea et al. 2024) 89.24(0.15) 91.02(0.10) 91.83(0.08) 85.50(0.15) 87.35(0.10) 87.84(0.08) 88.67(0.15) 90.21(0.10) 90.72(0.08)
+Physics-Encoded GNN(Saleh et al. 2024) 88.55(0.14) 90.44(0.09) 91.58(0.12) 84.81(0.14) 86.77(0.09) 87.59(0.12) 87.98(0.14) 89.63(0.09) 90.47(0.12)
+Transformer-KAN(Jiang et al. 2025) 90.89(0.12) 92.01(0.08) 93.22(0.07) 87.15(0.12) 88.34(0.08) 89.23(0.07) 90.32(0.12) 91.20(0.08) 92.11(0.07)
+Diffeomorphic Deformation(Gong et al. 2024) 91.14(0.10) 92.88(0.11) 94.33(0.09) 87.40(0.10) 89.21(0.11) 90.34(0.09) 90.57(0.10) 92.07(0.11) 93.22(0.09)
+PointNet(Qi et al. 2016) 94.23(0.11) 94.66(0.09) 96.54(0.09) 89.79(0.07) 90.11(0.07) 91.22(0.07) 94.00(0.08) 94.12(0.07) 95.01(0.06)
+U-Net(Mendizabal, Márquez-Neila, and Cotin 2020) 92.11(0.12) 93.56(0.10) 94.04(0.11) 88.23(0.08) 89.67(0.09) 90.01(0.08) 91.20(0.09) 92.33(0.10) 92.88(0.08)
+GraphGPS(Rampášek et al. 2022) 88.60(0.21) 89.45(0.15) 89.61(0.23) 87.33(0.06) 87.92(0.02) 88.01(0.04) 88.35(0.13) 88.91(0.10) 89.45(0.09)
+SAB(Lee et al. 2019) 93.42(0.11) 94.60(0.04) 96.03(0.04) 88.05(0.09) 89.91(0.11) 91.02(0.04) 92.54(0.06) 93.66(0.07) 94.42(0.06)
+Kelvinlet 65.24(0) 64.33(0) 64.82(0)
 Table 2: Performance of various models on linear simulations with individual graspers, combined graspers, and multitask learning.
-DCM (%) is reported for single-task and multi-task settings with standard deviations.
-Model Indiv. Graspers (DCM) Comb. Graspers (DCM) Multitask (DCM)
+DCM (%) is reported with standard deviations.
+“Base” denotes the original model, “+Res.” augments it with a Kelvinlet-based residual, and “+Reg.” applies a Kelvinlet-inspired regularizer.
+Model Indiv. (DCM) Comb. (DCM) Multi (DCM)
 Base +Res. +Reg. Base +Res. +Reg. Base +Res. +Reg.
-U -Net[22] 92.11 (0.12) 93.56 (0.10) 94.04 (0.11) 88.23 (0.08) 89.67 (0.09) 90.01 (0.08) 91.20 (0.09) 92.33 (0.10) 92.88 (0.08)
-GraphGPS[29] 88.60 (0.21) 89.45 (0.15) 89.61 (0.23) 87.33 (0.06) 87.92 (0.02) 88.01 (0.04) 88.35 (0.13) 88.91 (0.10) 89.45 (0.09)
-PointNet[28] 94.23 (0.11) 94.66 (0.09) 96.54 (0.09) 89.79 (0.07) 90.11 (0.07) 91.22 (0.07) 94.00 (0.08) 94.12 (0.07) 95.01 (0.06)
-SAB[19] 93.42 (0.11) 94.60 (0.04) 96.03 (0.04) 88.05 (0.09) 89.91 (0.11) 91.02 (0.04) 92.54 (0.06) 93.66 (0.07) 94.42 (0.06)
-Kelvinlet 65.24 (0) 64.33 (0) 64.82 (0)
+PhysGNN(Salehi and Giannacopoulos 2022) 78.05(0.13) 82.19(0.11) 82.79(0.10) 75.30(0.13) 79.18(0.11) 80.83(0.10) 77.73(0.13) 80.58(0.11) 81.34(0.10)
+SeeSaw(Docea et al. 2024) 77.17(0.15) 81.44(0.10) 81.97(0.08) 74.42(0.15) 78.43(0.10) 80.01(0.08) 76.85(0.15) 79.83(0.10) 80.52(0.08)
+Physics-Encoded GNN(Saleh et al. 2024) 76.48(0.14) 80.86(0.09) 81.72(0.12) 73.73(0.14) 77.85(0.09) 79.76(0.12) 76.16(0.14) 79.25(0.09) 80.27(0.12)
+Transformer-KAN(Jiang et al. 2025) 78.82(0.12) 82.43(0.08) 83.36(0.07) 76.07(0.12) 79.42(0.08) 81.40(0.07) 78.50(0.12) 80.82(0.08) 81.91(0.07)
+Diffeomorphic Deformation(Gong et al. 2024) 79.07(0.10) 83.30(0.11) 84.47(0.09) 76.32(0.10) 80.29(0.11) 82.51(0.09) 78.75(0.10) 81.69(0.11) 83.02(0.09)
+PointNet(Qi et al. 2016) 81.25(0.22) 84.65(0.19) 85.05(0.15) 79.56(0.18) 82.66(0.16) 83.33(0.11) 83.25(0.17) 84.75(0.13) 85.21(0.09)
+U-Net(Mendizabal, Márquez-Neila, and Cotin 2020) 80.56(0.10) 84.22(0.09) 84.88(0.11) 76.91(0.05) 79.83(0.08) 82.10(0.21) 78.10(0.12) 80.45(0.09) 81.33(0.14)
+GraphGPS(Rampášek et al. 2022) 77.52(0.17) 80.63(0.13) 82.11(0.09) 73.72(0.15) 77.92(0.10) 80.66(0.16) 76.24(0.18) 79.31(0.15) 80.42(0.10)
+SAB(Lee et al. 2019) 80.75(0.19) 84.44(0.07) 84.75(0.10) 78.88(0.11) 81.50(0.15) 82.83(0.16) 81.22(0.10) 83.01(0.08) 84.00(0.07)
+Kelvinlet 58.88(0) 51.67(0) 55.56(0)
 Table 3: Performance of various models on nonlinear simulations with individual graspers, combined graspers, and multitask learning.
-DCM (%) is reported for single-task and multi-task settings with standard deviations.
-Model Indiv. Graspers (DCM) Comb. Graspers (DCM) Multitask (DCM)
-Base +Res. +Reg. Base +Res. +Reg. Base +Res. +Reg.
-U -Net[22] 80.56 (0.10) 84.22 (0.09) 84.88 (0.11) 76.91 (0.05) 79.83 (0.08) 82.10 (0.21) 78.10 (0.12) 80.45 (0.09) 81.33 (0.14)
-GraphGPS[29] 77.52 (0.17) 80.63 (0.13) 82.11 (0.09) 73.72 (0.15) 77.92 (0.10) 80.66 (0.16) 76.24 (0.18) 79.31 (0.15) 80.42 (0.10)
-PointNet[28] 81.25 (0.22) 84.65 (0.19) 85.05 (0.15) 79.56 (0.18) 82.66 (0.16) 83.33 (0.11) 83.25 (0.17) 84.75 (0.13) 85.21 (0.09)
-SAB[19] 80.75 (0.19) 84.44 (0.07) 84.75 (0.10) 78.88 (0.11) 81.50 (0.15) 82.83 (0.16) 81.22 (0.10) 83.01 (0.08) 84.00 (0.07)
-Kelvinlet 58.88 (0) 51.67 (0) 55.56 (0)
+DCM (%) is reported with standard deviations.
+“Base” denotes the original model, “+Res.” augments it with a Kelvinlet-based residual, and “+Reg.” applies a Kelvinlet-inspired regularizer.
 To quantitatively assess model performance, we utilize the Deformation Capture Mean (DCM) metric, which measures how well a predicted deformation field approximates the ground-truth displacement field.
 The DCM is computed as:
-$\text{DCM}=100\times\left(1-\mathbb{E}_{(\mathbf{x}_{s},\mathbf{u}_{s})\sim p}\left[\frac{\left\|\mathbf{u}_{\text{true}}-\hat{\mathbf{u}}\right\|_{L_{2}(\Omega)}}{\|\mathbf{u}_{\text{true}}\|_{L_{2}(\Omega)}}\right]\right).$ (13)
-Table 1 shows that integrating Kelvinlet-based priors improves deformation accuracy across all models.
-Residual learning yields an average gain of +1.23%, while Kelvinlet regularization achieves a stronger improvement of +2.39%, with the highest gain of +3.19% observed in DiffoMorphic Deformation.
-PointNet benefits the most in absolute performance, reaching 96.54% DCM.
-The analytical Kelvinlet alone performs poorly (65.24%), underscoring the value of learning residuals.
-As seen in Table 2, our methods add minimal runtime overhead.
-Regularized variants match the base inference time, while residual versions increase latency by only $\sim$0.1ms.
-Kelvinlet inference alone is extremely fast (0.15ms) but not accurate enough, justifying its role as a prior.
-Overall, our approach enhances accuracy without compromising efficiency.
-Tables 3 and 4 report DCMs across models and settings.
-Each experiment runs with three random seeds, reporting means and standard deviations.
-Models perform better on linear cases, as nonlinear deformations introduce gravity effects, non-proportional stress-strain relationships, and greater sensitivity to perturbations.
-Notably, Kelvinlet priors consistently improve DCMs, despite being derived for linear elasticity, showing their effectiveness even in nonlinear regimes.
-Figure 4 illustrates the trade-off between inference time and accuracy for different models across linear and nonlinear simulations.
-The results indicate that Kelvinlet regularization consistently improves accuracy compared to the base models, as evidenced by the upward shift in solid markers relative to their transparent counterparts.
-However, this enhancement in accuracy does not increase computational cost, maintaining inference times within a comparable range.
-Additionally, models trained on multi-grasper data (black-edged markers) exhibit similar trends to their single-grasper counterparts, suggesting that Kelvinlet regularization generalizes well across different grasping conditions.
-All computations were performed on a single NVIDIA A6000 GPU, ensuring a consistent hardware environment for runtime measurements.
-5 Conclusion
-We introduce Neural-Augmented Kelvinlets, a physics-informed approach for real-time soft tissue deformation modeling that enhances neural network predictions in accuracy, physical consistency, and generalization while maintaining real-time performance.
-We utilize Kelvinlets in two settings: as a residual learning framework to refine predictions beyond an analytical baseline and as a regularization term to enforce physically plausible deformations.
-Large-scale FEM simulations provide a benchmark for evaluating data-driven deformation models, demonstrating that Kelvinlet integration significantly improves performance, particularly in complex multi-grasper scenarios.
-Our GPU-accelerated PyTorch implementation enables real-time inference, making it well-suited for surgical robotics and medical training, while a high-fidelity rendering pipeline ensures realistic soft tissue simulations in real time.
+$\text{DCM}=100\times\left(1-\mathbb{E}_{(\mathbf{x}_{s},\mathbf{u}_{s})\sim p}\left[\frac{\left\|\mathbf{u}_{\text{true}}-\hat{\mathbf{u}}\right\|_{L_{2}(\Omega)}}{\|\mathbf{u}_{\text{true}}\|_{L_{2}(\Omega)}}\right]\right)$ (13)
+Table 2 shows that Kelvinlet priors consistently improve deformation accuracy in the linear regime.
+Across all architectures, residual learning yields roughly 1 to 2 percentage points of DCM improvement, while Kelvinlet regularization adds a further 1 to 2 points, with the largest gain of +3.19% achieved by the diffeomorphic deformation network.
+PointNet achieves the highest absolute performance, reaching 96.54% DCM in the single grasper setting.
+In contrast, the analytical Kelvinlet alone remains substantially less accurate at 65.24% DCM, highlighting the importance of learned residual corrections.
+As shown in Table 1, the proposed methods incur negligible runtime overhead.
+Regularized models match the base inference time, while residual variants increase latency by only about 0.1ms.
+Although standalone Kelvinlet inference is extremely fast at 0.15ms, its accuracy is insufficient, supporting its role as a prior rather than a predictor.
+Overall, accuracy gains are achieved without sacrificing efficiency.
+Figure 4 illustrates the accuracy versus runtime trade off across linear and nonlinear simulations.
+Kelvinlet regularization consistently improves accuracy relative to the base models, as indicated by the upward shift of solid markers, while maintaining comparable inference times.
+Models trained on multi grasper data follow the same trend, indicating robust generalization across grasping conditions.
+All runtime measurements were obtained on a single NVIDIA A6000 GPU.
+Conclusion
+We present Neural Augmented Kelvinlets, a physics informed framework for real time soft tissue deformation modeling that improves accuracy, physical plausibility, and generalization while preserving real time performance.
+Kelvinlets are incorporated both as a residual correction mechanism and as a regularization prior that enforces physically consistent deformations.
+Extensive FEM simulations demonstrate clear performance gains, particularly in multi grasper scenarios.
+Our GPU accelerated PyTorch implementation enables real time inference and realistic visualization, making the method suitable for surgical robotics and medical training applications.
 Acknowledgments
-This project is supported by the Wellcome Leap Surgery: Assess/Validate/Expand (SAVE) program.
+This work was fully supported by the Wellcome-Leap SAVE program.